### v1

# Variable Selection in Convex Piecewise Linear Regression

Abstract
This paper presents Sparse Gradient Descent as a solution for variable selection in convex piecewise linear regression where the model is given as the maximum of $k$-affine functions $\bm{x}\mapsto\max_{j\in[k]}\langle\bm{a}_{j}^{\star},\bm{x}\rangle+b_{j}^{\star}$ for $j=1,\dots,k$. Here, $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ and $\{b_{j}^{\star}\}_{j=1}^{k}$ denote the ground-truth weight vectors and intercepts. A non-asymptotic local convergence analysis is provided for Sp-GD under sub-Gaussian noise when the covariate distribution satisfies sub-Gaussianity and anti-concentration property. When the model order and parameters are fixed, Sp-GD provides an $\epsilon$-accurate estimate given $\mathcal{O}(\max(\epsilon^{-2}\sigma_{z}^{2},1)s\log(d/s))$ observations where $\sigma_{z}^{2}$ denotes the noise variance. This also implies the exact parameter recovery by Sp-GD from $\mathcal{O}(s\log(d/s))$ noise-free observations. Since optimizing the squared loss for sparse max-affine is non-convex, an initialization scheme is proposed to provide a suitable initial estimate within the basin of attraction for Sp-GD, i.e. sufficiently accurate to invoke the convergence guarantees. The initialization scheme uses sparse principal component analysis to estimate the subspace spanned by $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ then applies an $r$-covering search to estimate the model parameters. A non-asymptotic analysis is presented for this initialization scheme when the covariates and noise samples follow Gaussian distributions. When the model order and parameters are fixed, this initialization scheme provides an $\epsilon$-accurate estimate given $\mathcal{O}(\epsilon^{-2}\max(\sigma_{z}^{4},\sigma_{z}^{2},1)s^{2}\log^{4}(d))$ observations. Numerical Monte Carlo results corroborate theoretical findings for Sp-GD and the initialization scheme.
Keywords— variable selection, non-linear regression, max-affine, piecewise linear
1 Introduction
We consider a multivariate regression problem where the target variable $y\in\mathbb{R}$ depends non-linearly on covariates in $\bm{x}\in\mathbb{R}^{d}$ and noise $z\in\mathbb{R}$ as
$y=f(\bm{x};\bm{\theta}^{\star})+z$ (1)
through a max-affine function $f:\mathbb{R}^{d}\to\mathbb{R}$ given by
$\bm{x}\in\mathbb{R}^{d}\mapsto f(\bm{x};\bm{\theta}^{\star})=\max_{1\leq j\leq k}\left(\langle\bm{a}^{\star}_{j},\bm{x}\rangle+b^{\star}_{j}\right),$ (2)
where $\bm{\theta}^{\star}\in\mathbb{R}^{k(d+1)}$ collects all ground-truth parameters $\{(\bm{a}_{j}^{\star},b_{j}^{\star})\}_{j=1}^{k}\subset\mathbb{R}^{d}\times\mathbb{R}$. The max-affine structure in (2) induces a class of convex piecewise linear functions. The set of max-affine functions provides an efficient approximation of a class of smooth convex functions [Balázs et al., 2015]. It is also considered a special instance of tropical algebra [Maragos et al., 2021]. Furthermore, we assume that only up to $s$ variables in $\bm{x}$ contribute to the evaluation of each linear model in (2), i.e.
$\left|\mathrm{supp}(\bm{a}^{\star}_{j})\right|\leq s,\quad\forall j=1,\dots,k,$ (3)
where $\mathrm{supp}(\cdot)$ denotes the index set for non-zero entries of the input vector. We refer to the estimation of the ground-truth parameters in $\bm{\theta}^{\star}$ from noisy samples $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ under the sparsity constraint in (3) as sparse max-affine regression. This is equivalent to convex piecewise linear regression with variable selection. The estimation procedure naturally implements variable selection as the support of the estimate identifies the active covariate variables.
The max-affine model in (2) proved useful in various applications such as computer vision and circuit design [Magnani and Boyd, 2009]. The target variable in these applications often has a non-linear dependence on a subset of the covariates. In high dimensional settings, i.e. the number of covariates is large, identifying the set of active covariates can significantly improve the generalization performance and the interpretability of the learned model [Wasserman and Roeder, 2009]. We illustrate these aspects over data fitting to an exponential polynomial, which extends the polynomial model so that the exponents are not necessarily integers. For example, in labor economics [Zhao et al., 2016, Eq. 15], the Gross Domestic Product (GDP), $G\in\mathbb{R}$, is written as a multivariate exponential polynomial
$G=c_{1}L^{\alpha}H^{\theta}S^{\gamma}D^{\delta}+c_{2}K+c_{3}SD/K+c_{4}$ (4)
with coefficients $\{c_{i}\}_{i=1}^{4}\subset\mathbb{R}$ and exponents $\{\alpha,\theta,\gamma,\delta\}\subset\mathbb{R}$, where $L,H,S,D,K$ respectively denote labor, human capital, innovation, investment, and capital stock. Employing a $\log$-$\log$ mapping, we can transform an exponential polynomial into a sparse max-affine function [Maragos et al., 2021, Boyd et al., 2004]. The procedure is described as follows: Let
$y=\varphi\,\mathrm{log}(G),\quad\bm{x}=\varphi\begin{bmatrix}\log L,\log H,\log S,\log D,\log K\end{bmatrix}^{\mathsf{T}}.$
for some $\varphi>0$. As $\varphi$ decreases toward 00, the expression of $y$ converges to a sparse max-affine function given by
$\underset{\varphi\rightarrow 0^{+}}{\mathrm{lim}}y=\underset{j\in\{1,2,3,4\}}{\mathrm{max}}\langle\bm{\theta}_{j},[\bm{x},1]\rangle,$
where
$[\bm{\theta}_{1},\ldots,\bm{\theta}_{4}]\triangleq\begin{bmatrix}\alpha&0&0&0\\ \theta&0&0&0\\ \gamma&0&1&0\\ \delta&0&1&0\\ 0&1&-1&0\\ 0&0&0&0\end{bmatrix}.$
More generally, even when the exact dependence like (4) is not known, sparse max-affine regression can identify active variables in each component of the target-covariate dependence modeled as an exponential polynomial.
Contributions:
In this paper, we present theoretical convergence guarantees for the sparse gradient descent algorithm (Sp-GD) that implements variable selection for max-affine regression. Since learning the regression parameters casts as a nonconvex optimization problem, it is important for Sp-GD to start from a suitable initialization. We propose an initialization scheme leveraging the sparsity structure backed by performance guarantees on its error bound. The theoretical results for both Sp-GD and the initialization scheme, presented respectively in pseudo Theorems 1.1 and 1.2, show that the sample complexity is governed by the number of active variables $s$ instead of the total number of variables $d$. Monte Carlo simulations in Sections 4.1 and 4.2 corroborate these theoretical guarantees.
Optimization algorithm:
Sp-GD is a variant of the projected gradient descent method. Sp-GD differs from standard projected gradient descent by utilizing a generalized gradient. This gradient is an extension to non-smooth functions, such as the piecewise linear functions we are working with [Hiriart-Urruty, 1979, Clarke, 1990]. Sp-GD moves along the generalized gradient with a set of adaptive step sizes and then projects to the feasible set of sparse parameters defined in (3). Since the nonlinear least squares problem for the model in (1) is non-convex, Sp-GD provides a local convergence guarantee presented in the following pseudo-theorem.
Theorem 1.1 (Informal).
Let the covariates and noise be sampled independently from Sub-Gaussian distributions. For fixed $k$ and ground-truth $\bm{\theta}^{\star}$ satisfying (3), with high proability, a suitably initialized Sp-GD converges linearly to an $\epsilon$-accurate estimate of $\bm{\theta}^{\star}$ given $\mathcal{O}\left(\max(\epsilon^{-2}\sigma_{z}^{2},1)s\log(d/s)\right)$ observations, where $\sigma_{z}^{2}$ denotes the noise variance.
This result can be compared to a line of research on plain max-affine regression without the sparsity constraint. Ghosh et al. [2021] presented non-asymptotic convergence analyses of the alternating minimization algorithm by Magnani and Boyd [2009] under random covariates and independent stochastic noise assumptions. Later, Kim and Lee [2024] showed that stochastic gradient descent provides comparable sample complexities and estimation errors with faster convergence.
Authors Algorithm Sample complexity
Ghosh et al. [2021] Alternating Minimization $\widetilde{\mathcal{O}}(d)$
Kim and Lee [2024] First Order Methods $\widetilde{\mathcal{O}}(d)$
This paper Sparse Gradient Descent $\widetilde{\mathcal{O}}\left(s\log({d}/{s})\right)$
Table 1: Comparison of estimation algorithms in the sample complexity for exact parameter recovery of max-affine models from noiseless observations.
Table 1 compares the required sample complexity for Sp-GD and previous parameter estimation algorithms for max-affine models. Sp-GD drops the sample complexity from $\mathcal{O}(d)$ required by previous algorithms to $\mathcal{O}(s\log(d/s))$ which is sub-linear. This marks a significant improvement especially when the number of active variables $s$ is significantly smaller than the total number of variables $d$. It is also worth noting that the convergence result in Theorem 1.1 applies to a practical implementation of SpGD. Specifically, the step sizes are given in an explicit form determined by the parameter estimates of the previous iteration. The step size for the $j$th block will concentrate around the inverse of the probability where the $j$th linear model in (2) achieves the maximum. In contrast, Kim and Lee [2024] used an unspecified constant step size to prove the local convergence of their first-order methods. Our step size strategy also applies to plain max-affine regression.
Initialization:
Recall that Theorem 1.1 requires a suitable initial estimate to guarantee local convergence of Sp-GD. To obtain this desired initialization, one may use the spectral method for max-affine regression [Ghosh et al., 2021]. Simply stated, their initialization scheme first employs principal component analysis (PCA) to estimate the span of $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$. Then a discrete search over an $r$-covering in the span of the principal components is applied. It requires $\mathcal{O}(\epsilon^{-2}d\log^{3}d)$ observations for their method to guarantee an $\epsilon$-accurate estimate with $r=\mathcal{O}(\epsilon)$ for fixed $\bm{\theta}^{\star}$ and noise level. This voids the gain due to Theorem 1.1. We propose a variant of the initialization by Ghosh et al. [2021] that substitutes PCA with sparse PCA (sPCA) and provides the desired initial estimate from fewer observations in the following special scenario: Suppose that the sparse coefficient vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ are simultaneously supported within a set of cardinality $s$, i.e.
$\left|\bigcup_{j=1}^{k}\mathrm{supp}(\bm{a}^{\star}_{j})\right|\leq s.$ (5)
Then the following pseudo-theorem quantifies the gain via the modified initialization scheme under this scenario.
Theorem 1.2 (Informal).
Let the covariates and noise be sampled independently from Gaussian distributions. Fix $k$ and ground-truth $\bm{\theta}^{\star}$ to satisfy the joint sparsity assumption in (5). Then, with high probability, the initialization via sPCA and $r$-covering search provides an $\epsilon$-accurate estimate of $\bm{\theta}^{\star}$ given $\mathcal{O}(\epsilon^{-2}\max(\sigma_{z}^{4},\sigma_{z}^{2},1)s^{2}\log^{4}d)$ observations and $r=\mathcal{O}(\epsilon)$ where $\sigma^{2}_{z}$ is the noise variance.
Compared to previous work [Ghosh et al., 2021], the initialization scheme leveraging the joint sparsity drops the linear dependence on the ambient dimension $d$ and replaces it with a quadratic dependence on $s$. Our implementation and theoretical analysis build on sPCA by semi-definite program by Vu et al. [2013] where the estimation error decays as $\mathcal{O}(\sqrt{s^{2}\log d/n})$. A recent framework improved the error bound to $\mathcal{O}(\sqrt{s\log d/n})$ under the assumption that data are obtained through a linear transform acting on special multivariate distributions [Wang et al., 2014]. However, this assumption is not satisfied by data generated with the max-affine model.
Related work:
Variable selection for linear regression models has been extensively studied in the literature (e.g. [George, 2000, Wasserman and Roeder, 2009]). Various optimization approaches including Lasso and basis pursuit have been studied [Hastie et al., 2009, Wainwright, 2019]. In practical applications, the covariate-target dependence often exhibits non-linearity not captured by linear models. There have been efforts to extend variable selection for linear regression to the non-linear models via basis change [Meier et al., 2009, Koltchinskii and Yuan, 2010]. They modeled the nonlinear function describing the target-covariate relation as a sparse superposition over a certain function basis (e.g. splines). Variable selection in their methods identifies active multivariate functions in the basis expansion. This approach is similar to the well-known order selection for polynomial regression [Peixoto, 1987]. Unlike the above model, similar to variable selection for linear regression, the sparse max-affine model directly identifies the active covariate variables to explain the target variable.
Variable selection has also been used along with dimensionality reduction for high-dimensional nonlinear models. The sparse minimum average variance estimator (sMAVE) [Wang and Yin, 2008] and sparse sliced inverse regression (sSIR) [Tan et al., 2020] considered the model where the target is given by a nonlinear function acting on linear combinations of a subset of covariates. Both sMAVE and sSIR estimate the span of the weight vectors in the linear combinations without attempting to learn the unknown non-linear function. Only asymptotic analyses of these algorithms exist in the high sample regime, i.e. the total number of samples goes to infinity while the model order remains fixed. The sparse max-affine model corresponds to a special case of their model where the non-linear function is known. Under this scenario, we provide a stronger non-asymptotic theoretical analysis for variable selection for the nonlinear max-affine model.
Paper organization:
Section 2 presents Sp-GD and its non-asymptotic theoretical guarantees as local analysis of sparse max-affine regression. Section 3.1 describes the initialization scheme the corresponding non-asymptotic theoretical guarantees. Section 4 presents the numerical results that corroborate the theoretical guarantees for Sp-GD and the initialization scheme. Section 5 provides final remarks and future directions.
Notation:
[…]
2 Local Analysis of Sparse Max-Affine Regression
2.1 Sparse Gradient Descent Algorithm
This section discusses the details of the Sp-GD algorithm. To simplify notation, we rewrite the max-affine model in (1) into a max-linear model
$y=\max_{j\in[k]}\langle\bm{\xi},\bm{\theta}_{j}^{\star}\rangle+z,$ (6)
where $\bm{\theta}_{j}^{\star}\triangleq[\bm{a}_{j}^{\star};b_{j}^{\star}]$ and $\bm{\xi}\triangleq[\bm{x};1]$ with the semicolon denoting vertical concatenation. Then the target sample $y_{i}$ is generated by (6) from the concatenated covariate sample $\bm{\xi}_{i}=[\bm{x}_{i};1]$ and noise sample $z_{i}$ for all $i\in[n]$. Let $\bm{\theta}^{\star}\triangleq[\bm{\theta}_{1}^{\star};\ldots;\bm{\theta}_{k}^{\star}]$ denote the vertical concatenation of all $k$ hyper-plane coefficient vectors $\{\bm{\theta}_{j}^{\star}\}_{j=1}^{k}\subset\mathbb{R}^{d+1}$. We consider an estimator of $\bm{\theta}^{\star}$ that minimizes the Mean Squared Error (MSE) loss function
$\ell\left([\bm{\theta}_{1};\ldots;\bm{\theta}_{k}]\right)\triangleq\frac{1}{2n}\sum_{i=1}^{n}\left(y_{i}-\max_{j\in[k]}\langle\bm{\xi}_{i},\bm{\theta}_{j}\rangle\right)^{2},$ (7)
under the constraint that all $\bm{\theta}_{j}$, for $j\in[k]$, belongs to $\Gamma_{s}$ defined by
$\Gamma_{s}\triangleq\left\{\bm{\varphi}\in\mathbb{R}^{(d+1)}:{\left\lVert[\bm{\varphi}]_{1:d}\right\rVert}_{0}\leq s,\right\},$ (8)
where ${\left\lVert\cdot\right\rVert}_{0}$ counts the number of nonzero entries and $[\bm{\varphi}]_{1:d}$ denotes the sub-vector of $\bm{\varphi}\in\mathbb{R}^{d+1}$ with the first $d$ entries.
Input: dataset $\{\bm{x}_{i},y_{i}\}_{i=1}^{n}$, sparsity level $s$, model rank $k$, and initial estimate $\bm{\theta}^{0}$
$t\leftarrow 0$
while stop condition is not satisfied do
       for $j\in\{1,\dots,k\}$ do
$\pi_{j}^{t}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\left\{\bm{x}_{i}\in\mathcal{C}_{j}(\bm{\theta}^{t})\right\}}$ (9)
if $\pi_{j}^{t}>0$ then
$\nabla_{\bm{\theta}_{j}}\ell(\bm{\theta}^{t})\leftarrow\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\left\{\bm{x}_{i}\in\mathcal{C}_{j}(\bm{\theta}^{t})\right\}}\left(\langle\bm{\xi}_{i},\bm{\theta}^{t}_{j}\rangle-y_{i}\right)\bm{\xi}_{i}$ (10)
$~{}~{}\bm{\alpha}_{j}^{t+1}\leftarrow\bm{\theta}_{j}^{t}-\left(\pi_{j}^{t}\right)^{-1}\nabla_{\bm{\theta}_{j}}{\ell}\left(\bm{\theta}^{t}\right)$ (11)
${}~{}\bm{\theta}_{j}^{t+1}\leftarrow\bm{\Psi}_{s}(\bm{\alpha}_{j}^{t+1})$
            else
                   ${}~{}\bm{\alpha}_{j}^{t+1}\leftarrow\bm{\theta}_{j}^{t}$
             end if
       end for
      $t\leftarrow t+1$
end while
Output: final estimate $\widehat{\bm{\theta}}\leftarrow\bm{\theta}^{t}$
Algorithm 1 Sparse Gradient Descent (Sp-GD)
Sp-GD is a variant of the projected gradient descent algorithm to pursue the above estimators, where the gradient is substituted by the generalized gradient [Hiriart-Urruty, 1979] and the step size varies across blocks adaptively with the iterates. We introduce a geometric object to describe the Sp-GD algorithm. Let
[…]
For any $\bm{\theta}\in\mathbb{R}^{k(d+1)}$, the open sets $\{\mathcal{C}_{j}(\bm{\theta})\}_{j=1}^{k}$ and their boundary in $\mathcal{V}(\bm{\theta})$ constructs a partition of $\mathbb{R}^{d}$ by satisfying
$\left(\bigcup\limits_{j=1}^{k}\mathcal{C}_{j}(\bm{\theta})\right)\cup\mathcal{V}(\bm{\theta})=\mathbb{R}^{d},\quad\left(\bigcup\limits_{j=1}^{k}\mathcal{C}_{j}(\bm{\theta})\right)\cap\mathcal{V}(y)=\varnothing,\quad\mathcal{C}_{j}(\bm{\theta})\cap\mathcal{C}_{l}(\bm{\theta})=\varnothing,\quad\forall l\neq j.$
The max-affine function in (6) is a special instance of tropical polynomials in the max-plus algebra [Maragos et al., 2021]. In this perspective, the sets $\{\mathcal{C}_{j}(\bm{\theta})\}_{j=1}^{k}$ and $\mathcal{V}(\bm{\theta})$ are called respectively as tropical open cells and tropical zero-set.
Algorithm 1 presents a pseudo code for the SpGD algorithm. Each iteration of the algorithm starts by updating the empirical probability $\pi_{j}^{t}$ that the covariates belong to then open set $\mathcal{C}_{j}(\bm{\theta}^{t})$ determined by the previous iterate $\bm{\theta}^{t}$ for all $j\in[k]$. Note that the evaluation of the indicator function in (9) can be saved and reused in the subsequent step in (10). If $\pi_{j}^{t}$ is non-zero, then we update the $j$th block by generalized gradient descent in (10) followed by the orthogonal projection to $\Gamma_{s}$ given by
$\bm{\Psi}_{s}(\bm{\alpha})=\mathop{\mathrm{argmin}}_{\widetilde{\bm{\alpha}}\in\Gamma_{s}}~{}{\left\lVert\bm{\alpha}-\widetilde{\bm{\alpha}}\right\rVert}_{2}^{2}.$
Otherwise, the $j$th block remains the same as the previous iterate. This update rule applies recursively until the algorithm converges by satisfying $\lVert\bm{\theta}^{t+1}-\bm{\theta}^{t}\rVert_{2}/\lVert\bm{\theta}^{t}\rVert_{2}$ less than a given threshold. Note that the step size for the gradient descent is adaptively evaluated as the reciprocal of $\pi_{j}^{t}$ which is determined solely by the previous iterate $\bm{\theta}^{t}$. In other words, this algorithm does not require tuning for the step size. In fact, the step size in (10) is always larger than $1$, which is quite different from the typical choices of step size for gradient descent (e.g. a small constant or a diminishing sequence). The next section presents theoretical guarantees of Sp-GD with this proposed step size. Furthermore, as shown in Section 4.1, this step size rule also makes Sp-GD empirically converge fast to the desired estimate.
2.2 Theoretical Analysis of Sp-GD
In this section, we present a local convergence analysis of Sp-GD under a set of covariate distributions determined by the following two properties.
Assumption 2.1 (Sub-Gaussianity).
The covariate vector $\bm{x}\in\mathbb{R}^{d}$ is zero-mean, isotropic, and $\eta$-sub-Gaussian, i.e. there exists $\eta>0$ such that
$\sup\limits_{\bm{u}\in\mathbb{S}^{d-1}}\mathbb{E}\left[\langle\bm{u},\bm{x}\rangle\right]\leq\exp\left({{\eta^{2}}/{2}}\right),$
where $\|\cdot\|_{\psi_{2}}$ and $\mathbb{S}^{d-1}$ denote the sub-Gaussian norm and the unit sphere in $\ell_{2}^{d}$, respectively.
Assumption 2.2 (Anti-Concentration).
There exist $\gamma,\zeta>0$ such that the covariate vector $\bm{x}\in\mathbb{R}^{d}$ satisfies
$\sup\limits_{\begin{subarray}{c}\bm{u}\in\mathbb{S}^{d-1},\lambda\in\mathbb{R}\end{subarray}}\mathbb{P}\left[\left(\langle\bm{u},\bm{x}\rangle+\lambda\right)^{2}\leq\epsilon\right]\leq\left(\gamma\epsilon\right)^{\zeta},\quad\forall\epsilon>0.$
We also introduce a set $\Theta(s,\kappa,\pi_{\min})$ that collects all $\bm{\theta}^{\star}\in\mathbb{R}^{k(d+1)}$ satisfying the following conditions: Recall that $\bm{\theta}^{\star}$ collects the weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}\subset\mathbb{R}^{d}$ and the bias terms $\{b_{j}^{\star}\}_{j=1}^{k}\in\mathbb{R}$ of the max-affine function (2) by $\bm{\theta}^{\star}=[[\bm{a}_{1}^{\star};b_{1}^{\star}];\ldots;[\bm{a}_{k}^{\star};b_{k}^{\star}]]$. First, the weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ satisfy the sparsity condition
${\left\lVert\bm{a}_{j}^{\star}\right\rVert}_{0}\leq s,\quad\forall j\in[k].$
Furthermore, any two distinct weight vectors are separated at least by a minimum discrepancy value $\kappa$, i.e.
$\min_{j^{\prime}\neq j}\|\bm{a}_{j}^{\star}-\bm{a}_{j^{\prime}}^{\star}\|_{2}\geq\kappa.$ (13)
Lastly, the probability that $j$th linear model achieves the maximum in (2) with random $\bm{x}$ should exceed a minimum threshold $\pi_{\min}>0$ for all $j\in[k]$, i.e.
$\min_{j\in[k]}\mathbb{P}\left(\bm{x}\in\mathcal{C}_{j}(\bm{\theta}^{\star})\right)\geq\pi_{\min}.$ (14)
Using the above definitions, we state the main result that shows the local convergence of Sp-GD in the following theorem.
Theorem 2.3.
Suppose that $\{(\bm{x}_{i},z_{i})\}_{i=1}^{n}$ are independent copies of a random vector $(\bm{x},z)$ where $\bm{x}\in\mathbb{R}^{d}$ and $z\in\mathbb{R}$ are independent, $\bm{x}$ satisfies Assumptions 2.12.2 with parameters $\eta,\gamma,\zeta>0$, and $z$ is zero-mean sub-Gaussian with variance $\sigma_{z}^{2}$. Then there exist absolute constants $C_{1},C_{2},R>0$, for which the following statement holds with probability at least $1-\delta$ for all $\bm{\theta}^{\star}\in\Theta(s,\kappa,\pi_{\min})$. If the initial estimate $\bm{\theta}^{0}$ belongs to a neighborhood of $\bm{\theta}^{\star}$ given by
$\mathcal{N}(\bm{\theta}^{\star},\kappa\rho):=\Bigg{\{}\bm{\theta}\in\mathbb{R}^{k(d+1)}\,:\,{\left\lVert\bm{\theta}-\bm{\theta}^{\star}\right\rVert}_{2}\leq\kappa\rho\Bigg{\}}$ (15)
with
$\rho:=\left[\frac{R\pi_{\min}^{3/4}}{4k^{2}}\cdot\log^{-1/2}\left(\frac{k^{2}}{R\pi_{\min}^{3/4}}\right)\right]\wedge\frac{1}{4}$ (16)
and
$n\geq C_{1}\left[s\log\left(\frac{n\vee d}{s}\right)+\log\left(\frac{k}{\delta}\right)\right]\left(\sigma_{z}^{2}\vee 1\vee\eta^{4}\right)k^{4}\pi_{\min}^{-4(1+\zeta^{-1})},$ (17)
then the sequence $\left(\bm{\theta}^{t}\right)_{t\in\mathbb{N}}$ generated by Sp-GD satisfies
$\left\|\bm{\theta}^{t}-\bm{\theta}^{\star}\right\|^{2}_{2}$ $\leq\tau^{t}\left\|\bm{\theta}^{0}-\bm{\theta}^{\star}\right\|^{2}_{2}+C_{2}\sigma^{2}_{z}\left(\frac{sk\log\left(n/s\right)+s\log\left({d}/{s}\right)+\log\left({1}/{\delta}\right)}{n}\right)$ (18)
for some $\tau\in(0,1)$ determined by $\pi_{\min}$, $k$, $\gamma$, $\zeta$ and $R$.
Remark 2.4.
A few remarks on the conditions and parameters in Theorem 2.3 are in order.
  1. 1.
    The exact form of constant $\tau\in(0,1)$ that determines the speed of convergence in (18) is provided as Equation (A) in Section A.
  2. 2.
    The parameter $\rho$ in (16) determines the size of the basin of attraction in (15) together with the minimum discrepancy value $\kappa$. In the “well-balanced” case, i.e. $\pi_{\min}=\Omega({1/k})$, we have that $\rho$ becomes $\Omega(k^{-11/4})$. Therefore, the basin of attraction shrinks only by the order $k$ of the max-affine model.
  3. 3.
    In the well-balanced case, if $\bm{x}\sim\mathrm{Normal}(\bm{0},\bm{I}_{d})$ (thus having $\zeta=1/2$, $\gamma=e$, and $\eta=1$), then the sample complexity requirement on $n$ becomes $\widetilde{\mathcal{O}}((\sigma_{z}^{2}\vee 1)sk^{8}\log d)$. Furthermore, for all $k\geq 2$, via an upper bound on $\tau$ obtained by evaluating (A) with $\rho=1/4$, we obtain that $\tau<10^{-3}$. Note that the expression of $\tau$ in (A) as a function of $\rho$ is monotone increasing.
Theorem 2.3 implies local linear convergence of Sp-GD in the noiseless case when the algorithm is properly initialized for the sub-Gaussian covariate model. The sample complexity scales linearly with $s$ significantly improving analogous results without variable selection [Ghosh et al., 2021, Kim and Lee, 2024]. Importantly, Sp-GD does not inflate the degree of dependence on the model order $k$ and dataset imbalance parameter $\pi_{\min}$ and maintains the same order $k^{4}\pi_{\min}^{-12}$ as plain GD and SGD [Kim and Lee, 2024]. Therefore, Sp-GD outperforms these algorithms even in the absence of sparsity in the max-affine model.
3 Initialization for Sparse Max-Affine Regression
3.1 Initialization Algorithms
The local convergence guarantees of Sp-GD provided in Theorem 2.3 require a suitable initial estimate. That is the case since the minimization of $\ell(\bm{\theta})$ in (7) under the sparsity constraint by (8) is non-convex, there exist multiple local minimizers, which hinders Sp-GD from converging to the global minimizer from an arbitrary initialization.
To bypass this issue, we compute an initial estimate in the basin of attraction of the ground truth so that the subsequent Sp-GD converges to the desired estimate. One may apply the initialization scheme for max-affine regression by Ghosh et al. [2021] while ignoring the sparsity constraint. However, their theoretical guarantees provide the desired accuracy when the sufficient number of observations scales at least linearly in the total number of variables $d$. Therefore, using this initialization scheme with Sp-GD yields a requirement on the sample complexity that is no longer dominated by the number of active variables $s$. To retain the gain due to Theorem 2.3, we propose an initialization scheme modified from the spectral initialization for max-affine regression [Ghosh et al., 2021] that provides an error rate depending on $s$ instead of $d$. To fulfill this objective, the initialization scheme additionally requires that the $s$-sparse weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ are jointly supported on a set with cardinality at most $s$. In other words, the ground-truth parameter vector $\bm{\theta}^{\star}=[[\bm{a}_{1}^{\star};b_{1}^{\star}];\ldots;[\bm{a}_{k}^{\star};b_{k}^{\star}]]$ belongs to the set defined by
$\Gamma_{\text{$s$-row-sparse}}\triangleq\left\{[\bm{\alpha}_{1};\dots;\bm{\alpha}_{k}]\in\mathbb{R}^{k(d+1)}:{\left\lVert\left(\sum_{j=1}^{k}[\bm{\alpha}_{j}]_{l}^{2}\right)_{l=1}^{d}\right\rVert}_{0}\leq s\right\},$ (19)
which is the set of all possible $\bm{\theta}^{\star}$ with jointly sparse weight vectors. The parameter initialization is a two-step process: (i) estimate the span of $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$, (ii) then estimate individual weight vectors $\{[\bm{a}_{j}^{\star};b_{j}^{\star}]\}_{j=1}^{k}$ from the estimated subspace.
Input: dataset $\{\bm{x}_{i},y_{i}\}_{i=1}^{n}$, sparsity level $s$, model order $k$, regularization parameter $\lambda$
$\bm{\widehat{m}}_{1}\leftarrow\sum_{i=1}^{n}y_{i}\bm{x}_{i},$ $\quad\bm{\widehat{M}}_{2}\leftarrow\sum_{i=1}^{n}y_{i}\left(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d}\right),$ $\quad\bm{\widehat{M}}\leftarrow\bm{\widehat{m}}_{1}\bm{\widehat{m}}_{1}^{\scriptscriptstyle{\textup{{T}}}}+\widehat{\bm{M}}_{2}$
$\bm{\widehat{P}}\leftarrow\displaystyle\mathop{\mathrm{argmax}}_{\widetilde{\bm{P}}\in\mathcal{F}^{k}_{d}}~{}\mathrm{tr}\left(\bm{\widehat{M}}^{\scriptscriptstyle{\textup{{T}}}}\widetilde{\bm{P}}\right)-\lambda\sum\limits_{i,j}\left|[\widetilde{\bm{P}}]_{i,j}\right|$triangleright where $\mathcal{F}^{k}_{d}$ is defined in (21)
$\mathcal{S}\leftarrow\left\{\text{indices of $s$-largest diagonal entries of $\widehat{\bm{P}}$}\right\},\quad\bm{\widehat{V}}\leftarrow\displaystyle\mathop{\operatorname*{argmin}}_{\widetilde{\bm{V}}^{\scriptscriptstyle{\textup{{T}}}}\widetilde{\bm{V}}=\bm{I}_{k}}{\left\lVert[\bm{\widehat{P}}]_{\mathcal{S},\mathcal{S}}-\widetilde{\bm{V}}\widetilde{\bm{V}}^{\scriptscriptstyle{\textup{{T}}}}\right\rVert}_{\mathrm{F}}$
Output: $k$-principle subspace estimate $\widehat{\bm{V}}\in\mathbb{R}^{s\times k}$, support set $\mathcal{S}$
Algorithm 2 Sparse Spectral Method for $k\leq s$
First, the subspace estimation algorithm is presented in Algorithm 2. This algorithm modifies upon the moment method for parameter estimation originally developed for mixtures of linear models [Chaganty and Liang, 2013, Zhang et al., 2014, Yi et al., 2016, Sedghi et al., 2016]. In Algorithm 2, $\widehat{\bm{m}}_{1}$ and $\widehat{\bm{M}}_{2}$ are respectively the first and second central moments weighted by the target values. In the non-sparse case, applying PCA to $\widehat{\bm{M}}$ and taking the first $k$-dominant eigenvectors yields a basis estimate of the subspace spanned by the ground-truth weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$. In our case, the joint sparsity of the dominant eigenvectors can be utilized to obtain a more accurate estimate of the subspace via sparse PCA (sPCA) [Zou et al., 2006]. Several convex relaxations were proposed to solve the sPCA problem [d’Aspremont et al., 2008, Amini and Wainwright, 2009, Zhang et al., 2012, Vu et al., 2013, Dey et al., 2018, Li and Xie, 2020]. Algorithm 2 uses the convex relaxation of sPCA by [Vu et al., 2013] that computes the $k$-dominant eigenvectors under the joint sparsity constraint. They formulated the estimation of sparse principal components as the following semidefinite program:
$\max_{\bm{P}\in\mathcal{F}^{k}_{d}}\mathrm{tr}\left(\widehat{\bm{M}}^{\top}\bm{P}\right)-\lambda\sum\limits_{i,j}\left|[{\bm{P}}]_{i,j}\right|$ (20)
for some positive constant $\lambda>0$, where $\mathcal{F}^{k}_{d}$ denotes the Fantope defined by
$\mathcal{F}^{k}_{d}\triangleq\left\{\bm{V}\bm{V}^{\scriptscriptstyle{\textup{{T}}}}:\bm{0}_{d}\preceq\bm{V}\bm{V}^{\scriptscriptstyle{\textup{{T}}}}\preceq\bm{I}_{d},~{}\mathrm{tr}\left(\bm{V}^{\scriptscriptstyle{\textup{{T}}}}\bm{V}\right)=k\right\}.$ (21)
Indeed, the Fantope $\mathcal{F}^{k}_{d}$ is obtained as the convex hull of all rank $k$-projection matrices [Overton and Womersley, 1992]. On one hand, the first term in the maximization in (20) measures the similarity between an element in the Fantope and the empirical moment matrix $\widehat{\bm{M}}$. On the other hand, the second term in (20) encourages sparsity by penalizing with the $\ell_{1}$ norm of all entries of the matrix. This relaxed version of sPCA is applied in the second step of Algorithm 2. The solution is calculated using the Alternating Direction Method of Multipliers (ADMM) algorithm presented in [Vu et al., 2013, Algorithm 1]. Since the result $\widehat{\bm{P}}$ is not necessarily a valid rank-$k$ projection matrix, we need post-processing in the following two steps. First, the algorithm recovers the support as $\mathcal{S}$ by thresholding the diagonal entries of $\widehat{\bm{P}}$. The final step in Algorithm 2 is to find the optimal projection of $[\bm{\widehat{P}}]_{\mathcal{S},\mathcal{S}}$ onto the set of all rank-$k$ projection matrices, which can be obtained by the $k$-dominant eigenvectors of $[\widehat{\bm{P}}]_{\mathcal{S},\mathcal{S}}$. Then $\widehat{\bm{V}}\in\mathbb{R}^{s\times k}$ will denote the Cholesky factor of the estimated rank-$k$ projection matrix.
Input: dataset $\{\bm{x}_{i},y_{i}\}_{i=1}^{n}$, model order $k$, subspace basis $\bm{\widehat{V}}$, separation $r\in(0,1)$, and support $\mathcal{S}$
$\mathcal{N}\leftarrow r$-covering of $B_{2}^{k+1}$
($\left(\{[\bm{w}_{j}^{\sharp};b_{j}^{\sharp}]\}_{j=1}^{k},c^{\sharp}\right)\leftarrow\mathop{\mathrm{argmin}}_{\{[\bm{w}_{j};b_{j}]\}_{j=1}^{k}\in\mathcal{N},c\geq 0}\frac{1}{n}\sum\limits_{i=1}^{n}\left(y_{i}-c\cdot\underset{j\in[k]}{\max}\langle[\bm{x}_{i}]_{\mathcal{S}}^{\mathrm{T}}\widehat{\bm{V}}w_{j}+b_{j}\rangle\right)^{2}$
for $j\in\{1,\dots,k\}$ do
       $[\bm{\theta}_{j}^{0}]_{\mathcal{S}}\leftarrow c^{\sharp}[\widehat{\bm{V}}\bm{w}_{j}^{\sharp};b_{j}^{\sharp}],\quad[\bm{\theta}_{j}^{0}]_{[d]\setminus\mathcal{S}}\leftarrow\bm{0}\in\mathbb{R}^{d-s}$
end for
Output: Initial model parameter estimate $\{\bm{\theta}_{j}^{0}\}_{j=1}^{k}$.
Algorithm 3 Discrete Search over Estimated Subspace [Ghosh et al., 2021]
Algorithm 2 identifies the joint support of $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ and estimates the subspace spanned by these vectors. This information is not sufficient to estimate the individual parameter vectors $\{[\bm{a}_{j}^{\star},b_{j}^{\star}]\}_{j=1}^{k}$. To approximate the parameter vectors up to a global scaling, Ghosh et al. [2021] proposed a discrete search over subsets of an $r$-covering $\mathcal{N}$ of the unit ball that satisfies $\min_{\bm{x}\in\mathcal{N}}\|\bm{x}-\bm{u}\|\leq r$ for all $\bm{u}\in B_{2}^{k+1}$ and $B_{2}^{k+1}\subset\bigcup_{\bm{w}\in\mathcal{N}}\left(\bm{w}+rB_{2}^{k+1}\right)$. The cost of constructing an $r$-covering and searching over it grows exponentially in the dimension. Due to the dimensionality reduction by Algorithm 2, the search dimension is the model order $k$, which is often much smaller than the ambient dimension $d$. If this is the case, then the discrete search is computationally feasible. This method also applies to the initialization of sparse max-affine regression. To make the manuscript self-contained, we summarize the discrete search algorithm by Ghosh et al. [2021] in Algorithm 3 using our notation.
3.2 Theoretical Analysis of Initialization
We present theoretical guarantees for Algorithm 2 in the following theorem.
Theorem 3.1.
Suppose that $k\leq s\leq d$. Let $\bm{x}\sim\mathrm{Normal}(\bm{0},\bm{I}_{d})$, $y$ be defined from $\bm{x}$ according to (6), and $\bm{P}\in\mathbb{R}^{d\times d}$ be the projection operator onto the span of $\{\bm{a}^{\star}_{j}\}_{j=1}^{k}$ that are jointly supported on $\mathcal{S}^{\star}$ with $|\mathcal{S}^{\star}|=s$. Then it holds with probability $1-n^{-11}$ that the estimates by Algorithm 2 satisfies $\mathcal{S}=\mathcal{S}^{\star}$ and
${\left\lVert\widehat{\bm{V}}\widehat{\bm{V}}^{\scriptscriptstyle{\textup{{T}}}}-[\bm{P}]_{\mathcal{S}^{\star},\mathcal{S}^{\star}}\right\rVert}_{\mathrm{F}}\leq Cs\cdot\frac{\varsigma^{2}+\sigma_{z}^{2}}{\delta_{\mathrm{gap}}}\cdot\left(\frac{\log^{2}(nd)}{n}\vee\frac{\log(nd)}{\sqrt{n}}\right)$ (22)
provided
$n\geq Cs^{2}\left(\frac{\varsigma^{2}\vee\sigma_{z}^{2}}{\delta_{\mathrm{gap}}}\vee\frac{\varsigma^{4}\vee\sigma_{z}^{4}}{\delta^{2}_{\mathrm{gap}}}\right)\left(\min_{j\in\mathcal{S}^{\star}}[\bm{P}]_{jj}\right)^{-2}\log^{2}(nd)$ (23)
for $\delta_{\mathrm{gap}}>0$, independent of $d$, where
$\varsigma\triangleq\max_{j\in[k]}\left({\left\lVert\bm{a}^{\star}_{j}\right\rVert}_{1}+|b_{j}^{\star}|\right).$ (24)
Ghosh et al. [2021] showed that their spectral initialization provides an $\epsilon$-accurate subspace estimation with $\widetilde{\mathcal{O}}(\epsilon^{-2}d)$ samples (when the ground-truth and hence the model order $k$ are fixed). In the sparse case of max-affine regression, Theorem 3.1 reduces the dependence of the sample complexity on $d$ from linear to logarithmic. Note that the spectral gap $\delta_{\mathrm{gap}}$ in Theorem 3.1 is also independent of $d$ similar to [Ghosh et al., 2021, Theorem 2].
Proof of Theorem 3.1.
Consider the population-level version of the empirical moment matrix $\widehat{\bm{M}}$ defined as
$\bm{M}=\bm{m}_{1}\bm{m}_{1}^{\scriptscriptstyle{\textup{{T}}}}+\bm{M}_{2},\quad\text{where}\quad\bm{m}_{1}=\mathbb{E}[y\bm{x}],\quad\text{and}\quad\bm{M}_{2}=\mathbb{E}\left[y(\bm{x}\bm{x}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d})\right].$ (25)
We will use the following known results about $\bm{M}$: i) The column space of $\bm{M}$ coincides with the $k$-dimensional subspace spanned by the ground-truth weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ [Ghosh et al., 2021, Lemma 3]; ii) There exists $\delta_{\mathrm{gap}}>0$, independent of $d$, such that the smallest nonzero eigenvalue of $\bm{M}$ is bounded from below by $\delta_{\mathrm{gap}}$ [Ghosh et al., 2021, Lemma 7]. Then [Vu et al., 2013, Theorem 3.1] provides a perturbation bound given by
$\|\widehat{\bm{P}}-\bm{P}\|_{\mathrm{F}}\leq\frac{4s}{\delta_{\mathrm{gap}}}\cdot{\left\lVert\bm{\widehat{M}}-\bm{M}\right\rVert}_{\infty}.$ (26)
Next, we derive an upper bound on $\|\widehat{\bm{M}}-\bm{M}\|_{\infty}$. The following lemma, whose proof is deferred to Appendix D, provides the concentration of $\widehat{\bm{M}}$ around $\bm{M}$.
Lemma 3.2.
Instate the assumptions in Theorem 3.1.We have that
$\mathbb{P}\left({\left\lVert\bm{\widehat{M}}-\bm{M}\right\rVert}_{\infty}\geq C(\varsigma^{2}+\sigma_{z}^{2})\left(\frac{\log^{2}(nd)}{n}\vee\frac{\log(nd)}{\sqrt{n}}\right)\right)\leq n^{-11}.$ (27)
By plugging Lemma 3.2 to (26), we obtain that it holds with probability $1-n^{-11}$ that
$\|\widehat{\bm{P}}-\bm{P}\|_{\mathrm{F}}\leq Cs\frac{\varsigma^{2}+\sigma_{z}^{2}}{\delta_{\mathrm{gap}}}\cdot\left(\frac{\log^{2}(nd)}{n}\vee\frac{\log(nd)}{\sqrt{n}}\right).$ (28)
We further proceed with the remainder of the proof under the event that (28) holds. The first assertion $\mathcal{S}=\mathcal{S}^{\star}$ follows from [Vu et al., 2013, Theorem 3.2] if
${\left\lVert\bm{\widehat{P}}-\bm{P}\right\rVert}_{\mathrm{F}}\leq\frac{1}{2}\cdot\min\limits_{\begin{subarray}{c}j\in\mathcal{S}^{\star}\end{subarray}}[\bm{P}]_{jj},$ (29)
which is satisfied by (23) and (28). Next, by the triangle inequality and the optimality of $\widehat{\bm{V}}$, we have
${\left\lVert\widehat{\bm{V}}\widehat{\bm{V}}^{\scriptscriptstyle{\textup{{T}}}}-[\bm{P}]_{\mathcal{S}^{\star},\mathcal{S}^{\star}}\right\rVert}_{\mathrm{F}}$ $\leq{\left\lVert\widehat{\bm{V}}\widehat{\bm{V}}^{\scriptscriptstyle{\textup{{T}}}}-[\bm{\widehat{P}}]_{\mathcal{S},\mathcal{S}}\right\rVert}_{\mathrm{F}}+{\left\lVert[\bm{\widehat{P}}]_{\mathcal{S},\mathcal{S}}-[\bm{P}]_{\mathcal{S}^{\star},\mathcal{S}^{\star}}\right\rVert}_{\mathrm{F}}$
$\leq 2\,{\left\lVert[\bm{\widehat{P}}]_{\mathcal{S},\mathcal{S}}-[\bm{P}]_{\mathcal{S}^{\star},\mathcal{S}^{\star}}\right\rVert}_{\mathrm{F}}=2\|\widehat{\bm{P}}-\bm{P}\|_{\mathrm{F}},$ (30)
where the last identity holds since we have shown $\mathcal{S}=\mathcal{S}^{\star}$. Combining (28) and (3.2) yields that the second assertion in (22) also follows from (23), which concludes the proof. ∎
Once the support $\mathcal{S}^{\star}$ is exactly recovered by Algorithm 2, the estimation accuracy of Algorithm 3 is provided by [Ghosh et al., 2021, Theorem 3] with $d$ substituted by $s$. We provide the statement of this result in our notation for completeness.
Theorem 3.3 (A paraphrase of [Ghosh et al., 2021, Theorem 3]).
Instate the assumptions in Theorem 3.1. Let $R_{\max}:=\max_{j\in[k]}{\left\lVert\bm{\theta}^{\star}_{j}\right\rVert}$. Then it holds with probability at least $1-n^{-11}$ that the initial parameter estimate $\{\bm{\theta}^{0}_{j}\}_{j=1}^{k}$ by Algorithm 3 with $\mathcal{S}=\mathcal{S}^{\star}$ satisfies
$\min_{\begin{subarray}{c}\pi\in\mathrm{perm}([k])\end{subarray}}\sum_{j=1}^{k}{\left\lVert\bm{\theta}^{\star}_{j}-\bm{\theta}^{0}_{\pi(j)}\right\rVert}^{2}\leq\frac{k^{4}}{\pi_{\min}^{3}}\left\{R_{\max}^{2}\left(r^{2}+{\left\lVert\widehat{\bm{V}}\widehat{\bm{V}}^{\scriptscriptstyle{\textup{{T}}}}-[\bm{P}]_{\mathcal{S}^{\star},\mathcal{S}^{\star}}\right\rVert}_{\mathrm{F}}^{2}\right)+\frac{\sigma^{2}_{z}\log(1+{1}/{r})}{n}\right\}$ (31)
provided that
$n\geq C\left(\frac{k}{\pi_{\min}}\right)^{4}\left\{\left[s\log\left({nk}\right)\log^{2}\left(\frac{k}{\pi_{\min}}\right)\right]\vee\left[\left(\frac{\sigma_{z}}{\Delta}\right)^{2}\log\left(1+\frac{1}{r}\right)\right]\right\},$ (32)
${\left\lVert\widehat{\bm{V}}\widehat{\bm{V}}^{\scriptscriptstyle{\textup{{T}}}}-[\bm{P}]_{\mathcal{S}^{\star},\mathcal{S}^{\star}}\right\rVert}^{2}_{\mathrm{F}}\leq\frac{\Delta^{4}\pi_{\min}^{3}}{64R_{\max}^{2}k^{4}},$ (33)
$r^{2}\leq\frac{\Delta^{4}\pi_{\min}^{5}}{64R_{\max}^{2}k^{6}\log(k\pi^{-1}_{\min})}.$ (34)
Remark 3.4.
The estimation error in Theorem 3.3 is stated using the minimum distance between the ground-truth parameters $\{\bm{\theta}^{\star}_{j}\}_{j=1}^{k}$ and all permutations of the estimates $\{\bm{\theta}^{0}_{j}\}_{j=1}^{k}$. On the other hand, Ghosh et al. [2021, Theorem 3] stated the estimation error as the minimum distance up to both permutation and scaling ambiguities. However, careful examination of their proof shows that their error bound applies to a particular scaling with the minimizer $c^{\sharp}$ in Algorithm 3. Therefore, [Ghosh et al., 2021, Theorem 3] also implies (31).
Finally, the performance of the entire initialization scheme using Algorithms 2 and 3 in succession is presented in the following theorem.
Theorem 3.5.
Instate the assumptions of Theorems 3.1 and 3.3. Let $\epsilon\in(0,1)$. Then it holds with probability at least $1-n^{-11}$ that applying Algorithms 2 and 3 in succession yields an initial estimate satisfying
$\min_{\begin{subarray}{c}\pi\in\mathrm{perm}([k])\end{subarray}}\left(\sum_{j=1}^{k}{\left\lVert\bm{\theta}^{\star}_{j}-\bm{\theta}^{0}_{\pi(j)}\right\rVert}^{2}\right)^{1/2}\leq\epsilon$ (35)
provided that
$n$ $\geq Cs^{2}\epsilon^{-2}\left(\frac{\varsigma^{2}\vee\sigma_{z}^{2}}{\delta_{\mathrm{gap}}}\vee\frac{\varsigma^{4}\vee\sigma_{z}^{4}}{\delta^{2}_{\mathrm{gap}}}\right)\left(\frac{R_{\max}k^{2}}{\Delta^{2}\pi_{\min}^{2}}\vee\frac{1}{\min\limits_{j\in[s]}[\bm{P}]_{jj}}\right)^{2}\log^{4}(nd)\log\left(1+\frac{1}{r}\right)$ (36)
and
$r^{2}\leq\frac{\Delta^{4}\pi_{\min}^{5}\epsilon^{2}}{32R_{\max}^{2}k^{6}\log(k\pi_{\min}^{-1})}.$ (37)
Proof.
The sample complexity condition (36) implies (23) and invokes Theorem 3.1 to satisfy $\mathcal{S}=\mathcal{S}^{\star}$ and (33). Furthermore, (36) and (37) respectively imply (32) and (34). Therefore, Theorem 3.3 is invoked to provide (31). Finally, (36) implies the upper bound in (31) is less than $\epsilon$, which is the assertion in (35) ∎
Theorem 3.5 reduces the sample complexity of the spectral initialization for max-affine regression by Ghosh et al. [2021] when the weight vectors satisfy the joint $s$-sparsity structure. Specifically, the linear dependence on the ambient dimension $d$ drops to polynomial dependence on $s$, significantly reducing the sample complexity when $s\ll d$. The order of the polynomial in $s$ depends on the geometry of the model parameters $\{\bm{\theta}^{\star}_{j}\}_{j=1}^{k}$. For example, if $\bm{1}\in\mathbb{R}^{s}$ is in the span of $\{\bm{\theta}^{\star}_{j}\}_{j=1}^{k}$, then it can be shown that $\min_{j\in[s]}[\bm{P}]_{jj}\geq 1/s$, whence the sample complexity in Theorem 3.5 becomes $\mathcal{O}(s^{4})$.
4 Numerical Results
4.1 Phase Transitions of Sp-GD
This subsection presents numerical results of the Sp-GD algorithm that corroborate the theoretical guarantees presented in Section 2.2. In the simulation, we initialized Sp-GD in two steps. First, we estimate the parameter subspace by Algorithm 2. Second, we apply a practical alternative to Algorithm 3 which randomly samples from the estimated subspace and chooses the best one that produces the smallest fit error after $10$ iterations of Sp-GD. We adopt this heuristic instead of Algorithm 3 for the following reasons. The first step of Algorithm 3 creates an $r$-covering of the unit $\ell_{2}$-ball $B^{k+1}$, where its cardinality scales as $\mathcal{O}(r^{-k})$ [Vershynin, 2018, Corollary 4.2.12]. Therefore, the exhaustive search in the second step over all elements in the $r$-covering becomes impractical as the parameter $r$ decreases. However, the accuracy of Algorithm 3 crucially depends on $r$. The estimation performance is evaluated via the median of the relative error between the true model coefficients $\bm{\theta}^{\star}\triangleq(\bm{\theta_{j}}^{\star})_{j=1}^{k}$ and the estimated coefficients $\widehat{\bm{\theta}}\triangleq(\widehat{\bm{\theta}}_{j})_{j=1}^{k}$ over $50$ Monte Carlo simulations. The relative error is defined via the optimal permutation of the affine model indices as
$\mathtt{err}(\widehat{\bm{\theta}})\triangleq\min_{\pi\in\mathrm{Perm}([k])}\log_{10}\left({\sum_{j=1}^{k}\|\widehat{\bm{\theta}}_{\pi(j)}-\bm{\theta}_{j}^{\star}\|_{2}^{2}/\sum_{j=1}^{k}\|\bm{\theta}_{j}^{\star}\|_{2}^{2}}\right),$
where $\mathrm{Perm}([k])$ denotes the set of all permutations on $[k]$.
Figure 1: Median of $\mathrm{err}(\widehat{\bm{\theta}})$ for different ($n$,$d$) pairs using 50 Monte Carlo iterations for $k=3$ and $s=25$ with Gaussian (top) and Uniform (bottom) covariate distributions in the noiseless case.
Fig. 1 shows the empirical phase transition by Sp-GD per the total number of covariates $d$ when the number of active covariates is fixed to $s=25$ and the model order is fixed to $k=3$ in the noiseless case. We observe the empirical phase transition for Gaussian and uniform distributions both of which satisfy the assumptions of Theorem 2.3. The phase transition occurs when $n$ scales as a logarithmic function of $d$, corroborating the sample complexity in Theorem 2.3. Next, Fig. 2 shows the empirical phase transition by Sp-GD per the number of active covariates $s$ when the total number of covariates and model order are fixed to $d=200$ and $k=3$, respectively. This figure corroborates that the sample complexity required to invoke the performance guarantee for Sp-GD scales sub-linearly in $s$ as $\mathcal{O}(s\log(d/s))$. We observe this scaling law when $s/k\geq 10$. On the other hand, when $s/k<10$ the transition boundary increases as $s$ decreases. This is based on the observation of the left edges of the plots in Fig. 2. The ground-truth parameters are randomly generated as independent and identically distributed with respect to the standard Gaussian distribution. In particular, the weight vectors are almost pairwise orthogonal when $s/k$ is sufficiently large (e.g. $s/k\geq 10$), which makes $\pi_{\min}\approx 1/k$. However, the correlations among the weight vectors increase as $s/k$ decreases, hence $\pi_{\min}$ decreases. This incurs the increase in the sample complexity in Theorem 2.3, which is aligned with the empirical observation on the phase transition boundary on the success regime, that is, $\mathtt{err}(\widehat{\bm{\theta}})\leq-2.5$. Finally, Fig. 3 shows the empirical phase transition by Sp-GD per the noise variance $\sigma_{z}^{2}$ when the total number of covariates, the number of active covariates, and the model order are fixed to $d=200$, $s=50$ and $k=3$, respectively. The empirical phase transition boundary is proportional to the noise variance once it exceeds a certain threshold. This corresponds to the multiplicative factor $\max(1,\sigma_{z}^{2})$ in (17).
Figure 2: Median of $\mathrm{err}(\widehat{\bm{\theta}})$ for different ($n$,$s$) pairs using 50 Monte Carlo iterations for $k=3$ and $d=200$ with Gaussian (left) and Uniform (right) covariate distributions. The red curves are fitted with respect to $s\log(d/s)$ at the phase transition boundary for both figures.
Figure 3: Median of $\mathrm{err}(\widehat{\bm{\theta}})$ for different ($n$,$\sigma_{z}^{2}$) pairs using 50 Monte Carlo iterations for $s=50$, $d=200$ and $k=3$ with Gaussian covariates and local initial estimate.
4.2 Subspace Estimation and Random Search for Initialization
This subsection delves into the detailed empirical analysis of the initialization method employed in the previous section. First, Fig. 4 demonstrates the gain of Algorithm 2 with SPCA over the analogous spectral method with PCA. The regularization parameter $\lambda>0$ required by Algorithm 2 is set by a parameter sweep over a sampling grid. Alternatively, this parameter can be tuned through cross-validation.
Figure 4: Projection error difference using PCA (red), and Algorithm 2 (blue), and dashed guidelines showing $1/\sqrt{n}$ decay (black) and $1/n$ decay (green) with $s=20$, $d=200$, $k=3$, $\sigma_{z}=0.1$ and 50 Monte Carlo iterations.
The subspace estimation error using Algorithm 2 shown in blue is significantly less than that by the PCA-based spectral method shown in red. Furthermore, the estimation error by Algorithm 2 decays at a rate between $1/\sqrt{n}$ and $1/n$. These observations are consistent with the error bound in Theorem 3.1.
Next, we investigate the empirical performance of the repeated random initialization that replaces the exhaustive discrete search in Algorithm 3 by random sampling followed by $10$ iterations. In particular, we observe the estimation accuracy as the function in the number of random initializations $M$. Fig. 5 compares the estimation error by the repeated random initialization when the parameter subspace is estimated by the PCA-based spectral initialization [Ghosh et al., 2021] (red) and by Algorithm 2 (blue). For both of the two subspace estimation methods, the initialization error monotonically decreases as the number of trials $M$ increases. However, the initialization performance by Algorithm 2 is more accurate since it provides better subspace estimation performance.
Figure 5: Parameter estimation error using PCA (red) and Algorithm 2 (blue) when followed by $M$ random samples with $s=20$, $d=200$, $k=3$, $\sigma_{z}=0.1$ averaged over 50 Monte Carlo iterations.
5 Discussion
We consider variable selection for a class of non-linear regression models given by the maximum of $k$ affine models $\bm{x}\mapsto\max_{j\in[k]}\langle\bm{a}_{j}^{\star},\bm{x}\rangle+b_{j}^{\star}$ for $j=1,\dots,k$ where $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ and $\{b_{j}^{\star}\}_{j=1}^{k}$ denote the ground-truth weight vectors and intercepts. The weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ satisfy the joint $s$-sparse structure as we assume that only $s$ out of $d$ covariates contributes to explaining the response variable. This paper proposes a variant of projected gradient algorithm called Sp-GD to estimate the sparse model parameters. We provide non-asymptotic local convergence guarantees for Sp-GD under independent sub-Gaussian noise when the covariates follow a sub-Gaussian distribution satisfying the anti-concentration property. Under these assumptions, when the ground-truth model order and parameters are fixed, a suitably initialized Sp-GD converges linearly to an $\epsilon$-accurate parameter estimate given $\mathcal{O}(\max(\epsilon^{-2}\sigma_{z}^{2},1)s\log(d/s))$. In particular when the observations are noise-free ($\sigma_{z}^{2}=0$), Sp-GD guarantees exact parameter recovery.
Since minimizing the squared loss of sparse max-affine models is non-convex, starting Sp-GD within the basin of attraction is crucial for its convergence to the desired estimate. For this purpose, we propose a modification of the spectral method by Ghosh et al. [2021] which estimates the span of the max-affine weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ so that the subspace estimation utilizes the jointly sparse structure in the weight vectors via sparse principal component analysis. Combined with the $r$-covering search over the estimated subspace, the initialization scheme provides an $\epsilon$-accurate estimate when $r=\mathcal{O}(\epsilon)$ given $\mathcal{O}(\epsilon^{-2}\max(\sigma_{z}^{4},\sigma_{z}^{2},1)s^{2}\log^{4}d)$ observations when the ground-truth model parameters are fixed, and the covariates and noise follow Gaussian distributions. The dominating factor of the sample complexity is $s^{2}$, which is significantly smaller than $d$ in the non-sparse case. One noteworthy limitation of the initialization is the $r$-covering search inherited from the previous work on non-sparse max-affine regression [Ghosh et al., 2021]. The cost of constructing the $r$-covering and the exhaustive search over it increases exponentially in the subspace dimension. Therefore, the $r$-covering search is not practical for small separation parameter $r$ which is necessary for accurate estimates. However, this is the only known algorithm with theoretical guarantees for max-affine parameter initialization. Therefore, it would be a fruitful future direction to develop a practical initialization scheme that avoids the $r$-covering search and provides theoretical performance guarantees. Alternatively, it would also be intriguing to explore potential analysis of Sp-GD from random initialization that will extend the known theoretical results on single-index models [Tan and Vershynin, 2019, Chandrasekher et al., 2022].
Appendix A Proof of Theorem 2.3
The proof is obtained by showing that each update in Sp-GD monotonically decreases the distance to the ground truth $\bm{\theta}^{\star}$ by a factor $\tau\in[0,1)$ up to an additive distortion and remains in the neighborhood of $\bm{\theta}^{\star}$, i.e.
$\left\|\bm{\theta}^{t+1}-\bm{\theta}^{\star}\right\|^{2}_{2}\leq\tau\left\|\bm{\theta}^{t}-\bm{\theta}^{\star}\right\|^{2}_{2}+\underbrace{C_{1}\sigma^{2}_{z}\left(\frac{sk\log\left(n/s\right)+s\log\left({d}/{s}\right)+\log\left({1}/{\delta}\right)}{n}\right)}_{\rho_{\mathrm{noise}}},$ (38a)
$\bm{\theta}^{t+1}\in\mathcal{N}(\bm{\theta}^{\star},\sqrt{2}\Delta\rho)$ (38b)
for some absolute constant $C_{1}$, hold for all $t\in\mathbb{N}\cup\{0\}$. We prove this statement by induction. Let $t$ be arbitrarily fixed. We assume that (38) holds for all previous iterates.
To show that (38a) also holds for the current iterate, we introduce the following notation. Let $\mathcal{S}^{\star}$ and $\mathcal{S}^{t}$ denote the joint support of $\{\bm{\theta}_{j}^{\star}\}_{j=1}^{k}$ and $\{\bm{\theta}_{j}^{t}\}_{j=1}^{k}$, respectively. The union of $\mathcal{S}^{\star}$ and $\mathcal{S}^{t}$ is denoted by $\varphi^{t}$. The step-size used in Sp-GD stated in Algorithm 1 is stated as
$\mu_{j}(\bm{\theta})=\left(\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\left\{\bm{x}_{i}\in\mathcal{C}_{j}(\bm{\theta})\right\}}\right)^{-1},\quad\forall j\in[k].$ (39)
For notational simplicity, let $\mu_{j}^{t}\triangleq\mu_{j}(\bm{\theta}^{t})$, $\mathcal{C}_{j}^{t}\triangleq\mathcal{C}_{j}(\bm{\theta}^{t})$, and $\mathcal{C}_{j}^{\star}\triangleq\mathcal{C}_{j}(\bm{\theta}^{\star})$ for all $j\in[k]$. Then the left-hand side of (38) is upper-bounded by
$\left\|\bm{\theta}^{t+1}-\bm{\theta}^{\star}\right\|_{2}$ $=\left\|\left(\bm{I}_{k}\otimes\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\right)\left(\bm{\theta}^{t+1}-\bm{\theta}^{\star}\right)\right\|_{2}$
$\leq\left\|\left(\bm{I}_{k}\otimes\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\right)\left(\bm{\theta}^{t+1}-\bm{\alpha}^{t+1}\right)\right\|_{2}+\left\|\left(\bm{I}_{k}\otimes\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\right)\left(\bm{\alpha}^{t+1}-\bm{\theta}^{\star}\right)\right\|_{2}$
$=\left(\sum_{j=1}^{k}\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{\theta}^{t+1}_{j}-\bm{\alpha}^{t+1}_{j}\right)\right\|_{2}^{2}\right)^{1/2}+\left(\sum_{j=1}^{k}\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{\alpha}^{t+1}_{j}-\bm{\theta}^{\star}_{j}\right)\right\|_{2}^{2}\right)^{1/2}$
$\leq 2\left(\sum_{j=1}^{k}\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{\alpha}^{t+1}_{j}-\bm{\theta}^{\star}_{j}\right)\right\|_{2}^{2}\right)^{1/2}=2\left(\sum_{j=1}^{k}\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{\theta}^{t}_{j}-\mu_{j}^{t}\nabla_{\bm{\theta}_{j}}\ell(\bm{\theta}^{t})-\bm{\theta}^{\star}_{j}\right)\right\|_{2}^{2}\right)^{1/2},$ (40)
where the second inequality holds since
$\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{\theta}^{t+1}_{j}-\bm{\alpha}_{j}^{t+1}\right)\right\|_{2}\leq\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{\alpha}_{j}^{t+1}-\bm{\theta}_{j}^{\star}\right)\right\|_{2},\quad\forall j\in[k]$
which follows from the fact that $\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\theta}_{j}^{t+1}=\bm{\theta}_{j}^{t+1}$ coincides with the projection of $\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\alpha}_{j}^{t+1}$ onto $\Gamma_{s}$, and $\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\theta}_{j}^{\star}=\bm{\theta}_{j}^{\star}$ belongs to $\Gamma_{s}$ for all $j\in[k]$.
Let $j\in[k]$ be arbitrarily fixed. We further proceed with the additional shorthand notations: $\bm{h}_{j}^{t}\triangleq\bm{\theta}_{j}^{t}-\bm{\theta_{j}}^{\star}$, $\bm{v}_{jj^{\prime}}^{t}\triangleq\bm{\theta}_{j}^{t}-\bm{\theta}_{j^{\prime}}^{t}$, and $\bm{v}^{\star}_{jj^{\prime}}\triangleq\bm{\theta}^{\star}_{j}-\bm{\theta}^{\star}_{j^{\prime}}$ for all $j^{\prime}\neq j\in[k]$. Then, by the definition of $\mathcal{S}^{\star}$, we have
$\bm{\widetilde{\Pi}}_{\mathcal{S}^{\star}}\bm{v}^{\star}_{jj^{\prime}}=\bm{\widetilde{\Pi}}_{\mathcal{S}^{\star}}\bm{\theta}^{\star}_{j}-\bm{\widetilde{\Pi}}_{\mathcal{S}^{\star}}\bm{\theta}^{\star}_{j^{\prime}}=\bm{\theta}^{\star}_{j}-\bm{\theta}^{\star}_{j^{\prime}}=\bm{v}^{\star}_{jj^{\prime}}$ (41)
and
$\bm{h}_{j}^{t}=\bm{\widetilde{\Pi}}_{\varphi^{t}}\bm{h}_{j}^{t}=\bm{\widetilde{\Pi}}_{\varphi^{t}\cap\varphi^{t+1}}\bm{h}_{j}^{t}+\bm{\widetilde{\Pi}}_{\varphi^{t}\backslash\varphi^{t+1}}\bm{h}_{j}^{t}=\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{h}_{j}^{t}+\bm{\widetilde{\Pi}}_{\varphi^{t}\backslash\varphi^{t+1}}\bm{h}_{j}^{t}.$ (42)
Recall that the partial gradient in the right-hand side of (A) is written as
$\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\nabla_{\bm{\theta}_{j}}\ell(\bm{\theta}^{t})=\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\left(\langle\bm{\xi}_{i},\bm{\theta}_{j}^{t}\rangle-\max_{j\in[k]}\langle\bm{\xi}_{i},\bm{\theta}_{j}^{\star}\rangle-z_{i}\right)\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\xi}_{i}.$ (43)
We can obtain the following decomposition
$\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\left(\langle\bm{\xi}_{i},\bm{\theta}_{j}^{t}\rangle-\max_{j\in[k]}\langle\bm{\xi}_{i},\bm{\theta}_{j}^{\star}\rangle\right)$
$=\sum_{j^{\prime}=1}^{k}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}^{\star}_{j^{\prime}}\}}\left(\langle\bm{\xi}_{i},\bm{\theta}_{j}^{t}-\bm{\theta}_{j}^{\star}+\bm{\theta}_{j}^{\star}-\bm{\theta}_{j^{\prime}}^{\star}\rangle\right)$
$=\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\langle\bm{\xi}_{i},\bm{h}_{j}^{t}\rangle+\sum_{j^{\prime}\neq j}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}^{\star}_{j^{\prime}}\}}\langle\bm{\xi}_{i},\bm{v}_{jj^{\prime}}^{\star}\rangle$
$=\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\langle\bm{\xi}_{i},\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{h}_{j}^{t}\rangle+\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\langle\bm{\xi}_{i},\bm{\widetilde{\Pi}}_{\varphi^{t}\backslash\varphi^{t+1}}\bm{h}_{j}^{t}\rangle+\sum_{j^{\prime}\neq j}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}^{\star}_{j^{\prime}}\}}\langle\bm{\xi}_{i},\bm{v}_{jj^{\prime}}^{\star}\rangle$ (44)
where the first equality follows from $\{\mathcal{C}^{\star}_{j^{\prime}}\}_{j^{\prime}=1}^{k}$ being a partition of $\mathbb{R}^{d}$, the second inequality follows from the definitions of $\bm{h}_{j}^{t}$ and $\bm{v}_{jj^{\prime}}^{\star}$, and the last equality follows from (42). We now use (A) to rewrite (43) as
$\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\nabla_{\bm{\theta}_{j}}\ell(\bm{\theta}^{t})$ $=\underbrace{\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\langle\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\xi}_{i},\bm{h}^{t}_{j}\rangle\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\xi}_{i}}_{\bm{p}_{j}}\quad+\underbrace{\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\langle\bm{\widetilde{\Pi}}_{\varphi^{t}\backslash\varphi^{t+1}}\bm{\xi}_{i},\bm{h}^{t}_{j}\rangle\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\xi}_{i}}_{\bm{q}_{j}}$ (45)
$\quad+\underbrace{\frac{1}{n}\sum_{\begin{subarray}{c}i=1\\ j^{\prime}\neq j\end{subarray}}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\langle\bm{\xi}_{i},\bm{v}_{jj^{\prime}}^{\star}\rangle\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\xi}_{i}}_{\bm{c}_{j}}\quad-\underbrace{\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}z_{i}\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\xi}_{i}}_{\bm{d}_{j}}.$
Plugging (45) into (A) yields
$\frac{1}{4}\left\|\bm{h}^{t+1}\right\|_{2}^{2}$ $\leq\sum_{j=1}^{k}\left[\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{h}^{t}_{j}-\mu_{j}^{t}\bm{p}_{j}\right)\right\|_{2}+\mu_{j}^{t}\left(\left\|\bm{q}_{j}\right\|_{2}+\left\|\bm{c}_{j}\right\|_{2}+\left\|\bm{d}_{j}\right\|_{2}\right)\right]^{2}.$ (46)
We now need to derive an upper bound on each term on the right-hand side of (46). Let $\pi_{j}^{t}\triangleq\mathbb{P}(\bm{x}\in\mathcal{C}_{j}(\bm{\theta}^{t}))$ and $\pi_{j}^{\star}\triangleq\mathbb{P}(\bm{x}\in\mathcal{C}_{j}(\bm{\theta}^{\star}))$. We now state a lemma which combines all the events that hold with high probability that are used for proving Theorem 2.3.
Lemma A.1.
Instate the assumptions and definitions in Theorem 2.3. Then, the following events hold jointly for all $t\in\mathbb{N}\cup\{0\}$ and $j\in[k]$ with probability at least $1-\delta$:
$\sup\limits_{\begin{subarray}{c}|\varphi|\leq s\end{subarray}}\left\|\bm{\widetilde{\Pi}}_{\varphi}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}^{t}\}}\left(\bm{\xi}_{i}\bm{\xi}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\varphi}\right\|\leq k^{-3/2}\pi_{\mathrm{min}}^{2(1+\zeta^{-1})},$ (47)
$\left|\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}^{t}\}}-\pi_{j}^{t}\right|\leq k^{-3/2}\pi_{\mathrm{min}}^{2(1+\zeta^{-1})},$ (48)
$\frac{1}{n}\sum_{j^{\prime}\neq j}\sum_{\begin{subarray}{c}i=1\end{subarray}}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}^{t}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\langle\bm{\widetilde{\Pi}}_{\mathcal{S}^{\star}}\bm{\xi}_{i},\bm{v}_{jj^{\prime}}^{\star}\rangle^{2}\leq\frac{2}{5\gamma k}\left(\frac{\pi_{\min}}{16}\right)^{1+\zeta^{-1}}\sum_{j^{\prime}\neq j}\|\bm{v}_{jj^{\prime}}^{t}-\bm{v}_{jj^{\prime}}^{\star}\|_{2}^{2},$ (49)
$\sup\limits_{\begin{subarray}{c}|\varphi|\leq s\end{subarray}}{\left\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}^{t}\}}z_{i}\bm{\widetilde{\Pi}}_{\varphi}[\bm{x}_{i};1]\right\rVert}_{2}\leq C\sigma_{z}\sqrt{\frac{sk\log\left({n}/{s}\right)+s\log\left({d}/{s}\right)+\log\left({1}/{\delta}\right)}{n}}\triangleq\tau_{\mathrm{noise}},$ (50)
$\sup\limits_{\begin{subarray}{c}|\varphi|\leq s\end{subarray}}{\left\lVert\frac{1}{\sqrt{n}}\bm{\widetilde{\Pi}}_{\varphi}[\bm{\xi}_{1},\ldots,\bm{\xi}_{n}]\right\rVert}\leq k^{-2}\pi_{\mathrm{min}}^{2(1+\zeta^{-1})}$ (51)
where $C>0$ is some absolute constant.
We proceed with the proof under the assumption that all the statements in Lemma A.1 hold. For notational simplicity, we will use the short hand $\epsilon_{\mathrm{min}}\triangleq k^{-3/2}\pi_{\mathrm{min}}^{2(1+\zeta^{-1})}$ repeatedly. We can now begin upper bounding the terms in (46). The first summand in (46) is upper-bounded as
$\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{h}^{t}_{j}-\mu\bm{p}_{j}\right)\|_{2}$
$\leq{\left\lVert\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{I}_{d+1}-\frac{\mu_{j}k}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}\}}\bm{\xi}_{i}\bm{\xi}_{i}^{\mathsf{T}}\right)\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\right\rVert}\cdot{\left\lVert\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{h}^{t}_{j}\right\rVert}_{2}.$ (52)
The first factor on the right-hand side can be upper bound by the triangle inequality as
${\left\lVert\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\left missing}(\bm{I}_{d+1}-\frac{\mu^{t}_{j}}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\bm{\xi}_{i}\bm{\xi}_{i}^{\mathsf{T}}\right)\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\right\rVert}$
$\leq\mu^{t}_{j}\left(\underbrace{\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\left(\bm{\xi}_{i}\bm{\xi}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\right\|}_{\mathcal{A}}+\underbrace{\left|\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}-\frac{1}{\mu^{t}_{j}}\right|}_{\mathcal{B}}\right).$ (53)
By (47) we have that $\mathcal{A}\leq\epsilon_{\mathrm{min}}$, $\mathcal{B}=0$ by the definition of $\mu_{j}$. Also, since $\bm{\theta}^{t}\in\mathcal{N}(\bm{\theta}^{\star},\sqrt{2}\Delta\rho)$, by Lemma B.3 we have that
$(1-\varrho)\leq\frac{\pi_{j}^{t}}{\pi_{j}^{\star}}\leq\left(\frac{1-\varrho}{1-2\varrho}\right),\quad\varrho\triangleq{R}^{2\zeta}{k^{-2(1+\zeta^{-1})}}$ (54)
Therefore, using (48) and (54), we can upper bound the step size as
$\mu^{t}_{j}\triangleq\frac{1}{\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}}\leq\frac{1}{\pi^{t}_{j}-\epsilon_{\mathrm{min}}}\leq\frac{1}{(1-\varrho)\pi^{\star}_{j}-\epsilon_{\mathrm{min}}}$ (55)
Finally, the first summand in (46) is upper-bounded as
$\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{h}^{t}_{j}-\mu_{j}^{t}\bm{p}_{j}\right)\|_{2}\leq\frac{\epsilon_{\mathrm{min}}}{(1-\varrho)\pi^{\star}_{j}-\epsilon_{\mathrm{min}}}{\left\lVert\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{h}^{t}_{j}\right\rVert}_{2}.$ (56)
We can use (47) again to bound the second summand of (46) as
$\mu^{t}_{j}\|\bm{q}_{j}\|_{2}\leq\mu^{t}_{j}\epsilon\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{h}^{t}_{j}\|_{2}\leq\frac{\epsilon_{\mathrm{min}}}{(1-\varrho)\pi^{\star}_{j}-\epsilon_{\mathrm{min}}}.$
The vector $\bm{c}_{j}$ in the last term of (45) is factorized as $\bm{c}_{j}=\bm{\widetilde{\Pi}}_{\varphi^{t+1}}[\bm{\xi}_{1},\ldots,\bm{\xi}_{n}]\bm{v}$, where
$\bm{v}\triangleq\sum_{j^{\prime}\neq j}\begin{bmatrix}\mathbb{1}_{\{\bm{x}_{1}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\langle\bm{\widetilde{\Pi}}_{\mathcal{S}^{\star}}\bm{\xi}_{1},\bm{v}_{jj^{\prime}}^{\star}\rangle\\ \vdots\\ \mathbb{1}_{\{\bm{x}_{n}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\langle\bm{\widetilde{\Pi}}_{\mathcal{S}^{\star}}\bm{\xi}_{n},\bm{v}_{jj^{\prime}}^{\star}\rangle\end{bmatrix}.$
Therefore, we have
$\|\bm{c}_{j}\|_{2}\leq{\left\lVert\frac{1}{\sqrt{n}}\bm{\widetilde{\Pi}}_{\varphi^{t+1}}[\bm{\xi}_{1},\ldots,\bm{\xi}_{n}]\right\rVert}\cdot{\left\lVert\frac{1}{\sqrt{n}}\bm{v}\right\rVert}_{2}\leq\frac{\epsilon_{\mathrm{min}}}{\sqrt{k}}{\left\lVert\frac{1}{\sqrt{n}}\bm{v}\right\rVert}_{2},$ (57)
where the second inequality follows from (51). Next we bound the last term in (57) as
$\frac{1}{n}{\left\lVert\bm{v}\right\rVert}^{2}_{2}$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{j^{\prime}:j^{\prime}\neq j}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\langle\bm{\widetilde{\Pi}}_{\mathcal{S}^{\star}}\bm{\xi}_{i},\bm{v}_{jj^{\prime}}^{\star}\rangle^{2}$
$\leq\underbrace{\frac{2}{5\gamma}\left(\frac{\pi_{\min}}{16}\right)^{1+\zeta^{-1}}k^{-1}}_{\lambda^{2}}\sum_{j^{\prime}:j^{\prime}\neq j}\|\bm{v}_{jj^{\prime}}^{t}-\bm{v}_{jj^{\prime}}^{\star}\|_{2}^{2}$
$=\lambda^{2}\sum_{j^{\prime}:j^{\prime}\neq j}\|\bm{h}_{j}^{t}-\bm{h}_{j^{\prime}}^{t}\|_{2}^{2}$
$\leq\lambda^{2}\sum_{j^{\prime}:j^{\prime}\neq j}\left(\|\bm{h}_{j}^{t}\|_{2}^{2}+\|\bm{h}_{j^{\prime}}^{t}\|_{2}^{2}\right),$ (58)
where the first inequality follows from (49). Since the $\ell_{1}$ norm dominates the $\ell_{2}$ norm, we can write (A) as
$\|\frac{1}{\sqrt{n}}\bm{v}\|_{2}\leq\lambda\sum_{j^{\prime}:j^{\prime}\neq j}\left(\|\bm{h}_{j}^{t}\|_{2}+\|\bm{h}_{j^{\prime}}^{t}\|_{2}\right).$ (59)
Now, it remains to bound the last term on the right-hand side of (46) using (50) such that
$\|\bm{d}\|_{2}\leq\tau_{\mathrm{noise}}.$ (60)
Finally plugging the above upper bounds into (46) yields
$\|\bm{h}^{t+1}\|^{2}_{2}$
$\leq 4\sum_{j=1}^{k}\Bigg{\{}\frac{\epsilon_{\mathrm{min}}}{(1-\varrho)\pi^{\star}_{j}-\epsilon_{\mathrm{min}}}\bigg{[}\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{h}^{t}_{j}\|_{2}+\|\bm{\widetilde{\Pi}}_{\varphi^{t}\backslash\varphi^{t+1}}\bm{h}^{t}_{j}\|_{2}+4\lambda(k-1)\frac{\epsilon_{\mathrm{min}}}{\sqrt{k}}\|\bm{h}^{t}_{j}\|_{2}+\tau_{\mathrm{noise}}/\epsilon_{\mathrm{min}}\bigg{]}\Bigg{\}}^{2}$
$\leq 4\left[\frac{\epsilon_{\mathrm{min}}}{(1-\varrho)\pi^{\star}_{j}-\epsilon_{\mathrm{min}}}\right]^{2}\sum_{j=1}^{k}\left\{\bigg{[}\sqrt{2}+4\lambda(k-1)\frac{\epsilon_{\mathrm{min}}}{\sqrt{k}}\bigg{]}\|\bm{h}^{t}_{j}\|_{2}+\tau_{\mathrm{noise}}/\epsilon_{\mathrm{min}}\right\}^{2}$
$\leq\underbrace{2\left\{\frac{\epsilon_{\mathrm{min}}}{(1-\varrho)\pi_{\mathrm{min}}-\epsilon_{\mathrm{min}}}\bigg{[}\sqrt{2}+4\lambda(k-1)\frac{\epsilon_{\mathrm{min}}}{\sqrt{k}})\bigg{]}\right\}^{2}}_{\tau}\sum_{j=1}^{k}\|\bm{h}^{t}_{j}\|^{2}_{2}+8k\left[\frac{\tau_{\mathrm{noise}}}{(1-\varrho)\pi_{\min}-\epsilon_{\mathrm{min}}}\right]^{2}$
$=\tau\|\bm{h}^{t}\|^{2}_{2}+C\sigma^{2}_{z}\left(\frac{sk\log\left(\frac{n}{s}\right)+s\log\left(\frac{d}{s}\right)+\log\left(\frac{1}{\delta}\right)}{n}\right),$ (61)
for $\tau\in[0,1)$. The second inequality follows from
$\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{h}^{t}_{j}\|_{2}+\|\bm{\widetilde{\Pi}}_{\varphi^{t}\backslash\varphi^{t+1}}\bm{h}^{t}_{j}\|_{2}\leq\sqrt{2}\|\bm{\widetilde{\Pi}}_{\varphi^{t}}\bm{h}^{t}_{j}\|_{2},$ (62)
and the third inequality follows trivially from $\pi_{j}^{\star}\geq\pi_{\mathrm{min}}^{\star}$ for all $j\in[k]$ which finally verifies the first assertion in (38a). By the recursive nature of (38a), we have that
$\left\|\bm{\theta}^{t+1}-\bm{\theta}^{\star}\right\|^{2}_{2}\leq\tau^{t+1}\left\|\bm{\theta}^{0}-\bm{\theta}^{\star}\right\|^{2}_{2}+\frac{\rho_{\mathrm{noise}}}{1-\tau}.$ (63)
This first term on the right-hand side of (63) is upper bounded as
$\tau^{t+1}\left\|\bm{\theta}^{0}-\bm{\theta}^{\star}\right\|^{2}_{2}\leq(\Delta\rho)^{2}.$ (64)
Next, the second term on the right-hand side of (63) is upper bounded as
$\frac{\rho_{\mathrm{noise}}}{1-\tau}\overset{\mathrm{(i)}}{\leq}\frac{\rho_{\mathrm{noise}}}{1-\tau}k^{-3}\pi_{\min}^{2(1+\zeta^{-1})}C_{1}/C\overset{\mathrm{(ii)}}{\leq}(\Delta\rho)^{2},$ (65)
where (i) follows from the sample complexity requirement in (17), and (ii) follows for large enough $C>0$. Combining (64) and (65) implies
$\left\|\bm{\theta}^{t+1}-\bm{\theta}^{\star}\right\|^{2}_{2}\leq 2(\Delta\rho)^{2}\implies\bm{\theta}^{t+1}\in\mathcal{N}(\bm{\theta}^{\star},\sqrt{2}\Delta\rho),$ (66)
which concludes the proof using the strong law of induction.
Appendix B Auxiliary Lemmas for Theorem 2.3
This section will introduce several lemmas used to prove Theorem 2.3. To state these lemmas, we provide the following definitions. First, define the collection of all possible support sets of cardinality $s$ as
$\mathcal{Z}_{s}\triangleq\left\{\varphi\subset[d]:|\varphi|=s\right\}.$
Next, the set of all polytopes determined by $k$ jointly $s$-sparse halfspaces is defined as
$\mathcal{P}_{k,d,s}=\bigcup_{\varphi\in\mathcal{Z}_{s}}\mathcal{P}_{k,d}\left(\varphi\right),$ (67)
where
$\mathcal{P}_{k,d}(\varphi)\triangleq\left\{\bm{x}\in\mathbb{R}^{d}:[\bm{x}]_{\varphi^{\mathsf{c}}}=\bm{0},~{}M[\bm{x}]_{\varphi}\geq\bm{b},~{}M\in\mathbb{R}^{k\times s},~{}\bm{b}\in\mathbb{R}^{k}\right\}.$ (68)
We use $\bm{\Pi}_{\varphi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to denote the orthogonal projection onto the subspace spanned by the standard basis vectors $\{\bm{e}_{j}\}_{j\in\varphi}$ for $\varphi\subset[d]$. Then the augmented operator $\bm{\widetilde{\Pi}}_{\varphi}:\mathbb{R}^{d+1}\rightarrow\mathbb{R}^{d+1}$ is defined as
$\bm{\widetilde{\Pi}}_{\varphi}=\left[{\begin{array}[]{cc}\bm{\Pi}_{\varphi}&\bm{0}_{d\times 1}\\ \bm{0}_{1\times d}&1\\ \end{array}}\right].$ (69)
We now proceed with the statement of the lemmas. The next lemma is used to upper bound the worst-case operator norm of sub-Gaussian matrices with independent and jointly sparse columns.
Lemma B.1.
Let $\{\bm{x}_{i}\}_{i=1}^{n}$ be independent copies of a random vector $\bm{x}\in\mathbb{R}^{d}$ which satisfies Assumptions 2.1 and 2.2. Let $\{\omega_{i}\}_{i=1}^{n}\in\{0,1\}^{n}$ be fixed with $\sum_{i=1}^{n}\omega_{i}=\ell>0$. Then for all $\epsilon\in[0,1]$, there exists an absolute constant $C>0$ where it holds with probability at least $1-\delta$ that
$\underset{\varphi\in\mathcal{Z}_{s}}{\mathrm{sup}}{\left\lVert\bm{\widetilde{\Pi}}_{\varphi}\sum_{i=1}^{n}\omega_{i}\left([\bm{x}_{i};1][\bm{x}_{i};1]^{{\scriptscriptstyle{\textup{{T}}}}}-\bm{I}_{d+1}\right)\bm{\widetilde{\Pi}}_{\varphi}\right\rVert}\leq\ell\epsilon$ (70)
if
$l\geq C(\eta\vee 1)^{4}\epsilon^{-2}\left[s\log\left(\frac{d}{s}\right)+\log\left(\frac{1}{\delta}\right)\right].$ (71)
Proof.
The proof of the lemma relies on the unitary invariance of the spectral norm. Let $\{r_{i}\}_{i=1}^{n}$ be independent copies of the random variable $r$ following the Rademacher distribution. Also, let $\bm{\widetilde{\xi}}_{i}\triangleq r_{i}[\bm{x}_{i};1]$ for all $i\in[n]$. Then for $\bm{u}\in\mathbb{R}^{d}$ and $\lambda\in\mathbb{R}$ and for all $i\in[n]$ we have
$\mathbb{E}\left[\exp\left(\langle[\bm{u};\lambda],\bm{\widetilde{\xi}}_{i}\rangle\right)\right]$ $=\frac{e^{\lambda}}{2}\mathbb{E}\left[\exp\left(\langle\bm{u},\bm{x}_{i}\rangle\right)\right]+\frac{e^{-\lambda}}{2}\mathbb{E}\left[\exp\left(-\langle\bm{u},\bm{x}_{i}\rangle\right)\right]$
$\overset{\mathrm{(i)}}{\leq}\frac{1}{2}\exp\left(\frac{{\left\lVert\bm{u}\right\rVert}^{2}\eta^{2}}{2}\right)\left(e^{\lambda}+e^{-\lambda}\right)$
$\overset{\mathrm{(i)}}{\leq}\exp\left(\frac{{\left\lVert\bm{u}\right\rVert}^{2}\eta^{2}+\lambda^{2}}{2}\right)\leq\exp\left(\frac{(\eta\vee 1)^{2}{\left\lVert[\bm{u};\lambda]\right\rVert}^{2}}{2}\right),$ (72)
where (i) follows since $\{\bm{x}_{i}\}_{i=1}^{n}$ are $\eta$-sub-Gaussian random vectors, and (ii) follows using $e^{a^{2}/2}\geq\left(e^{a}+e^{-a}\right)/2$ for all $a\in\mathbb{R}$. Therefore, we have that $\{\bm{\widetilde{\xi}}\}_{i=1}^{n}$ are identical copies of a random vector $\bm{\widetilde{\xi}}$ which satisfies Assumption 2.1 and is $(\eta\vee 1)$-sub-Gaussian. The proof of this lemma becomes a direct application of the union bound by inflating the probability of error by $|\mathcal{Z}_{s}|=$ $\binom{d}{s}$ $\leq\left(\frac{ed}{s}\right)^{s}$ to the statement in [Wainwright, 2019, Theorem 6.5]. ∎
The next lemma presents that the empirical measure of sparse polytopes concentrates around the expectation.
Lemma B.2.
Let $\{\bm{x}_{i}\}_{i=1}^{n}$ be independent copies of a random vector $\bm{x}$ satisfying Assumption 2.1. Then there exists an absolute constant $C$ for which it holds with probability at least $1-\delta$ that
$\sup_{\begin{subarray}{c}\mathcal{C}\in\mathcal{P}_{k,d,s}\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}\}}-\mathbb{P}(\bm{x}\in\mathcal{C})\right|\leq\epsilon,$ (73)
if
$n\geq C\epsilon^{-2}\left[sk\log\left(\frac{n}{s}\right)+s\log\left(\frac{d}{s}\right)+\log\left(\frac{1}{\delta}\right)\right].$ (74)
Proof of Lemma B.2.
The proof of this lemma is obtained by a direct application of the union bound to the statement in [Kim and Lee, 2024, Corollary 6.7]. The result of [Kim and Lee, 2024, Corollary 6.7] provides the concentration inequality in (73) when the supremum is over $\mathcal{P}_{k,d}(\varphi)$ for a fixed $\varphi$ instead of the union over $\varphi\in\mathcal{Z}_{s}$. To apply the union bound argument, we inflate the probability of error by the factor $|\mathcal{Z}_{s}|=$ $\binom{d}{s}$ $\leq\left(\frac{ed}{s}\right)^{s}$. ∎
To bound the closeness of the empirical measure in Lemma B.2 to the expectation with ground truth model parameters, we present the following lemma.
Lemma B.3.
Suppose that $\bm{x}\in\mathbb{R}^{d}$ satisfies Assumptions 2.1 and 2.2. Let $\{\mathcal{C}_{j}(\bm{\theta})\}_{j=1}^{k}$ and $\{\mathcal{C}_{j}(\bm{\theta}^{\star})\}_{j=1}^{k}$ be respectively defined by $\bm{\theta}$ and $\bm{\theta}^{\star}\in\mathbb{R}^{dk}$ according to (12). Let $\mathcal{N}(\bm{\theta}^{\star},\Delta\rho)$ be defined with $R>0$ as in (15) and $\varrho\triangleq CR^{2\zeta}k^{-2(1+\zeta^{-1})}$. If $\bm{\theta}\in\mathcal{N}(\bm{\theta}^{\star},\Delta\rho)$, then we have
$(1-\varrho)\leq\frac{\mathbb{P}(\bm{x}\in\mathcal{C}_{j}(\bm{\theta}))}{\mathbb{P}(\bm{x}\in\mathcal{C}_{j}(\bm{\theta}^{\star}))}\leq\left(\frac{1-\varrho}{1-2\varrho}\right),\quad\forall j\in[k].$ (75)
The proof of this lemma is deferred to the Appendix D.3. The next lemma shows a tail bound on the sparsity-constrained operator norm of partial sums of the centered outer products of covariates.
Lemma B.4.
Let $\{\bm{x}_{i}\}_{i=1}^{n}$ be independent copies of a random vector $\bm{x}$ which satisfies Assumptions 2.1 and 2.2. Then there exists an absolute constant $C$ for which it holds with probability at least $1-\delta$ that
$\sup\limits_{\begin{subarray}{c}\mathcal{C}\in\mathcal{P}_{k,d,s}\\ \varphi\in\mathcal{Z}_{s}\end{subarray}}\left\|\bm{\widetilde{\Pi}}_{\varphi}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}\}}\left([\bm{x}_{i};1][\bm{x}_{i};1]^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\varphi}\right\|\leq\epsilon$ (76)
if
$n\geq C(\eta\vee 1)^{4}\epsilon^{-2}\left[sk\log\left(\frac{n}{s}\right)+s\log\left(\frac{d}{s}\right)+\log\left(\frac{1}{\delta}\right)\right].$ (77)
The proof of this lemma is deferred to Appendix D.4 The statement of the next lemma will require the notation
$v_{jj^{\prime}}\triangleq\bm{\theta}_{j}-\bm{\theta}_{j^{\prime}},\quad v^{\star}_{jj^{\prime}}\triangleq\bm{\theta}^{\star}_{j}-\bm{\theta}^{\star}_{j^{\prime}},\quad\forall j^{\prime}\neq j\in[k].$ (78)
Lemma B.5.
Let $\{\mathcal{C}_{j}(\bm{\theta})\}_{j=1}^{k}$ and $\{\mathcal{C}_{j}(\bm{\theta}^{\star})\}_{j=1}^{k}$ be respectively defined by $\bm{\theta}$ and $\bm{\theta}^{\star}\in\mathbb{R}^{dk}$ according to (12). Assume that $\bm{x}\in\mathbb{R}^{d}$ satisfies Assumptions 2.2 and 2.1. Fix $\delta\in(0,1/e)$ and $R>0$. Assume $\bm{\theta}\in\mathcal{N}(\bm{\theta}^{\star},\Delta\rho)$ as defined in (15) with $\rho$ as defined in (16). Then there exists an absolute constant $C$ for which it holds with probability at least $1-\delta$ that
$\frac{1}{n}\sum_{j^{\prime}\neq j}\sum_{\begin{subarray}{c}i=1\end{subarray}}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}(\bm{\theta})\cap\mathcal{C}_{j^{\prime}}(\bm{\theta}^{\star})\}}\langle\bm{\xi}_{i},\bm{v}_{jj^{\prime}}^{\star}\rangle^{2}\leq\frac{2}{5\gamma k}\left(\frac{\pi_{\min}}{16}\right)^{1+\zeta^{-1}}\sum_{j^{\prime}\neq j}\|\bm{v}_{jj^{\prime}}-\bm{v}_{jj^{\prime}}^{\star}\|_{2}^{2}$ (79)
if
$n\geq Ck^{4}\pi_{\mathrm{min}}^{-4(1+\zeta^{-1})}\left[s\log\left(\frac{n\vee d}{s}\right)+\log\left(\frac{k}{\delta}\right)\right].$ (80)
Proof.
The proof of this lemma is a direct application of the union bound by inflating the probability of error by $|\mathcal{Z}_{s}|=$$\binom{d}{s}$ $\leq\left(\frac{ed}{s}\right)^{s}$ to the statement in [Kim and Lee, 2024, Lemma 7.7]. ∎
The next lemma provides an upper bound of a noise-related term that will appear in the proof of the main theorem.
Lemma B.6.
Let $\{\bm{x}_{i}\}_{i=1}^{n}$ be independent copies of a random vector $\bm{x}$ satisfying Assumptions 2.1 and 2.2. Let $\{z_{i}\}_{i=1}^{n}$ be i.i.d. sub-Gaussian random variables with zero mean and variance $\sigma^{2}_{z}$, independent of everything else. Then there exists an absolute constant $C$ for which it holds with probability at least $1-\delta$ that
$\sup\limits_{\begin{subarray}{c}\mathcal{C}\in\mathcal{P}_{k,d,s}\\ \varphi\in\mathcal{Z}_{s}\end{subarray}}{\left\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}\}}z_{i}\bm{\widetilde{\Pi}}_{\varphi}[\bm{x}_{i};1]\right\rVert}_{2}\leq C\sigma_{z}\sqrt{\frac{sk\log\left(n/s\right)+s\log\left(d/s\right)+\log\left({1}/{\delta}\right)}{n}}.$ (81)
Proof.
The proof of lemma follows directly from applying the union bound to the statement in [Kim and Lee, 2024, lemma 8.1, Eq. 46]. Since that statement only considers the supremum over non-sparse polytopes, whereas we consider the supremum over jointly sparse polytopes and sparse vector supports, we must inflate the probability of error $\delta$ by $|\mathcal{Z}_{s}|^{2}=$$\binom{d}{s}^{2}$ $\leq\left(\frac{ed}{s}\right)^{2s}$. ∎
Appendix C Auxiliary Lemmas for Theorem 3.5
The first lemma aims to show that the empirical moment matrix $\widehat{\bm{M}}$, defined in 2, is not rank deficient. In other words, its first $k$ dominant eigenvectors are a basis for the span of $\{\bm{\theta}_{j}^{\star}\}_{j=1}^{k}$. To state this lemma, we will make use of the definitions in (25) and (24). The first lemma stated next is borrowed from [Ghosh et al., 2021].
Lemma C.1.
[Ghosh et al., 2021, Lemma 7] Let $\bm{x}\sim\mathrm{Normal}(\bm{0},\bm{I}_{d})$ , and $y$ be defined from $\bm{x}$ according to (6). Also, assume that $k\leq d$. Then we have that the combination of the first and second moments satisfies
$\bm{M}\succeq\bm{0}_{d},\quad\mathrm{rank}(\bm{M})=k,\quad\lambda_{k}(\bm{M})\geq\delta_{\mathrm{gap}},$
for some numerical constant $\delta_{\mathrm{gap}}>0$ independent of the ambient dimension $d$.
The next lemma states the concentration of the empirical moments around their expectations.
Lemma C.2.
Let $\bm{x}\sim\mathrm{Normal}(\bm{0},\bm{I}_{d})$ , and $y$ be defined from $\bm{x}$ according to (6). Then there exists absolute constants $C_{1},C_{2}>0$ such that
$\mathbb{P}\left({\left\lVert\bm{m}_{1}\bm{-}\bm{\widehat{m}_{1}}\right\rVert}_{\infty}\geq C_{1}(\varsigma+\sigma_{z})\frac{\log(nd)}{\sqrt{n}}\right)\leq n^{-11},$
$\mathbb{P}\left({\left\lVert\bm{M}_{2}-\bm{\widehat{M}}_{2}\right\rVert}_{\infty}\geq C_{2}(\varsigma+\sigma_{z})\log(nd)\left(\frac{1}{\sqrt{n}}\vee\frac{\sqrt[6]{\log(nd)}}{\sqrt[3]{n^{2}}}\right)\right)\leq n^{-12}.$ (82)
A direct consequence of these two lemmas is the concentration of moment matrix $\widehat{\bm{M}}$ around its expectation $\bm{M}$ stated in Lemma 3.2.
Lemma C.3 ( Maximum of Sub-exponential Random Variables).
Let $\{X_{j}\}_{j=1}^{k}$ have ${\left\lVert X_{j}\right\rVert}_{\psi_{1}}<\infty$ for all $j\in[k]$, then we have that
${\left\lVert\max_{j\in[k]}|X_{j}|\right\rVert}_{\psi_{1}}\leq\log_{2}(2k)\max_{j\in[k]}{\left\lVert X_{j}\right\rVert}_{\psi_{1}}.$
Proof.
The proof of this lemma follows from
$\mathbb{P}(\max_{j\in[k]}|X_{j}|>t)$ $\leq\sum_{j=1}^{k}\mathbb{P}(|X_{j}|>t)\leq\min\left(1,\sum_{j=1}^{k}2\exp\left\{{\frac{-t}{C{\left\lVert X_{j}\right\rVert}}_{\psi_{1}}}\right\}\right),$
$\leq\min\left(1,2k\exp{\frac{-t}{C\underset{j\in[k]}{\max}{\left\lVert X_{j}\right\rVert}_{\psi_{1}}}}\right)=\min\left(1,\exp\left\{\frac{-t}{C\underset{j\in[k]}{\max}{\left\lVert X_{j}\right\rVert}_{\psi_{1}}}+\log 2k\right\}\right),$
$\leq\min\left(1,2\exp{\frac{-t}{C\log_{2}(2k)\underset{j\in[k]}{\max}{\left\lVert X_{j}\right\rVert}_{\psi_{1}}}}\right),$
for all $t>0$ where the second inequality follows from the tail probability of any sub-exponential random variable [Vershynin, 2018, Proposition 2.7.1]. Again, comparing this tail probability with the last inequality yields the assertion in this lemma. A simple graphical example is provided in Figure 6, assuming that $C\underset{j\in[k]}{\max}{\left\lVert X_{j}\right\rVert}_{\psi_{1}}=1$, where we compare the last two tail probabilities in the statement of the proof. ∎
Figure 6: Comparison of tail probabilities as a function of the threshold $t>0$ showing the trivial bound (blue), the tail $2\exp\{\frac{-t}{\log_{2}(2k)}\}$ (green), and $\exp\{-t+\log(2k)\}$ (red) with $k=2$.
Appendix D Proof of Lemmas
D.1 Proof of Lemma 3.2
The proof of this lemma follows from
${\left\lVert\bm{\widehat{M}}-\bm{M}\right\rVert}_{\infty}$ $\leq{\left\lVert\bm{\widehat{M}}_{2}-\bm{M}_{2}\right\rVert}_{\infty}+{\left\lVert\bm{\widehat{m}}_{1}\bm{\widehat{m}}_{1}^{\scriptscriptstyle{\textup{{T}}}}-\bm{m}_{1}\bm{m}_{1}^{\scriptscriptstyle{\textup{{T}}}}\right\rVert}_{\infty}$
$\leq{\left\lVert\bm{\widehat{M}}_{2}-\bm{M}_{2}\right\rVert}_{\infty}+{\left\lVert\bm{\widehat{m}}_{1}-\bm{m}_{1}\right\rVert}_{\infty}^{2}+2{\left\lVert\bm{\widehat{m}}_{1}-\bm{m}_{1}\right\rVert}_{\infty}\cdot{\left\lVert\bm{m}_{1}\right\rVert}_{\infty}$
$\leq C(\varsigma^{2}+\sigma_{z}^{2})\left(\frac{\log^{2}(nd)}{n}\vee\frac{\log(nd)}{\sqrt{n}}\right),$ (83)
where the last inequality follows from both Lemma C.1 and C.2.
D.2 Proof of Lemma C.2
Recall that a sub-Gaussian random vector $\bm{x}\in\mathbb{R}^{d}$ satisfies
$\mathbb{P}\left({\left\lVert\bm{x}\right\rVert}_{\infty}\geq t\right)\leq 2d\exp\left\{-\frac{ct^{2}}{{\left\lVert\bm{x}\right\rVert}^{2}_{\psi_{2}}}\right\},$ (84)
for some absolute constant $c>0$. Next, define two events as
$\mathcal{E}_{i}=\left\{{\left\lVert\bm{x}_{i}\right\rVert}_{\infty}\leq 5\sqrt{\log(end)}\right\},\quad\mathcal{F}_{i}=\left\{z_{i}\leq 5\sigma_{z}\sqrt{\log(n)}\right\},\quad\forall i\in[n],$ (85)
where each event holds with probability at least $1-C_{1}n^{-12}$ for some absolute constant $C_{1}>0$. Next, for all $i\in[n]$, define the truncated covariate vector $\bm{\widetilde{x}}_{i}=\bm{x}_{i}\mathbb{1}_{\mathcal{E}_{i}}$. With this definition we can begin bounding the moment concentrations. We notice that $\bm{\widetilde{x}}_{i}-\bm{x}_{i}=\bm{\widetilde{x}}_{i}-(\mathbb{1}_{\{\mathcal{E}_{i}\}}+\mathbb{1}_{\{\mathcal{E}^{\mathsf{c}}_{i}\}})\bm{x}_{i}=-\mathbb{1}_{\{\mathcal{E}^{\mathsf{c}}_{i}\}}\bm{x}_{i}$, and $\mathbb{1}_{\mathcal{E}_{i}}\max\limits_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\widetilde{\bm{x}}_{i},1]\rangle=\mathbb{1}_{\mathcal{E}_{i}}\max\limits_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i},1]\rangle$ for all $i\in[n]$. Therefore, the difference of the first moments can be written as
$\bm{\widehat{m}}_{1}-\bm{m}_{1}$ $=\frac{1}{n}\sum\limits_{i=1}^{n}\left(\bm{x}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x};1]\rangle-\mathbb{E}\left[\bm{x}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x};1]\rangle\right]+z_{i}\bm{x}_{i}\right)$
$=\frac{1}{n}\sum\limits_{i=1}^{n}\left(\bm{x}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x};1]\rangle-\bm{\widetilde{x}}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{\widetilde{x}}_{i};1]\rangle+\mathbb{E}\left[\bm{\widetilde{x}}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{\widetilde{x}}_{i};1]\rangle\right]-\mathbb{E}\left[\bm{x}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle\right]\right)$
$\quad+\frac{1}{n}\sum\limits_{i=1}^{n}\left(\bm{\widetilde{x}}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{\widetilde{x}}_{i};1]\rangle-\mathbb{E}\left[\bm{\widetilde{x}}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{\widetilde{x}}_{i};1]\rangle\right]\right)+\frac{1}{n}\sum\limits_{i=1}^{n}z_{i}\bm{x}_{i}$
$=\underbrace{\frac{1}{n}\sum\limits_{i=1}^{n}\left(\mathbb{1}_{\mathcal{E}^{\mathsf{c}}_{i}}\bm{x}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle-\mathbb{E}\left[\mathbb{1}_{\mathcal{E}^{\mathsf{c}}_{i}}\bm{x}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle\right]\right)}_{\mathcal{M}_{1}}$
$\quad+\underbrace{\frac{1}{n}\sum\limits_{i=1}^{n}\left(\bm{\widetilde{x}}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{\widetilde{x}}_{i};1]\rangle-\mathbb{E}\left[\bm{\widetilde{x}}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{\widetilde{x}}_{i};1]\rangle\right]\right)}_{\mathcal{M}_{2}}+\frac{1}{n}\sum\limits_{i=1}^{n}z_{i}\bm{x}_{i}$ (86)
We now bound the max norm of the quantities in (D.2). First, we prove that each summand in $\mathcal{M}_{1}$ is sub-exponential. For this purpose, we write
${\left\lVert\mathbb{1}_{\mathcal{E}^{\mathsf{c}}_{i}}\bm{x}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle\right\rVert}_{\psi_{1}}$ $\leq{\left\lVert\mathbb{1}_{\mathcal{E}^{\mathsf{c}}_{i}}\bm{x}_{i}\right\rVert}_{\psi_{2}}{\left\lVert\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle\right\rVert}_{\psi_{2}}$
$\overset{\mathrm{(i)}}{\leq}C{\left\lVert\bm{x}_{i}\right\rVert}_{\psi_{2}}\cdot\sqrt{\log(k)}\max_{j\in[k]}{\left\lVert\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle\right\rVert}_{\psi_{2}}$
$\leq C\sqrt{\log(k)}\max_{j\in[k]}\left({\left\lVert\bm{a}_{j}^{\star}\right\rVert}_{2}+|b_{j}^{\star}|\right)\overset{\mathrm{(ii)}}{\leq}C\varsigma\sqrt{\log(k)},$ (87)
where (i) follows from Lemma C.3 and the fact that the indicator function is always dominated by $1$; and (ii) follows by recalling the definition $\varsigma=\underset{j\in[k]}{\max}({\left\lVert\bm{a}^{\star}_{j}\right\rVert}_{1}+|b_{j}^{\star}|)$ from (24) and the fact the $\ell_{1}$ norm dominates the $\ell_{2}$ norm. Therefore, we can apply Bernstein’s inequality for sub-exponential random variables along with the union bound over $d$ entries to get
$\mathbb{P}\left({\left\lVert\mathcal{M}_{1}\right\rVert}_{\infty}\geq C\epsilon\sqrt{\frac{\log(k)}{n}}\right)\leq de^{-\epsilon}.$ (88)
Choosing $\epsilon=12\frac{\log(nd)}{\log(k)}$, we get
$\mathbb{P}\left({\left\lVert\mathcal{M}_{1}\right\rVert}_{\infty}\geq C\frac{\log(nd)}{\sqrt{n}}\right)\leq n^{-12}.$ (89)
Next, we show that each summand in $\mathcal{M}_{2}$ is sub-Gaussian. For this purpose, we present a bound as
$\left|\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{\widetilde{x}}_{i};1]\rangle\right|$ $\leq\max_{j\in[k]}\left[\left|\langle\bm{a}_{j}^{\star},\bm{\widetilde{x}_{i}}\rangle\right|+|b_{j}^{\star}|\right]\leq\max_{j\in[k]}\left[{\left\lVert\bm{a}_{j}^{\star}\right\rVert}_{1}\cdot{\left\lVert\bm{\widetilde{x}_{i}}\right\rVert}_{\infty}+|b_{j}^{\star}|\right]$
$\leq\max_{j\in[k]}\left[5{\left\lVert\bm{a}^{\star}_{j}\right\rVert}_{1}\sqrt{\log(end)}+|b_{j}^{\star}|\right]\leq 5\varsigma\sqrt{\log(end)},\quad\forall i\in[n],$ (90)
using the definition of $\varsigma=\underset{j\in[k]}{\max}({\left\lVert\bm{a}^{\star}_{j}\right\rVert}_{1}+|b_{j}^{\star}|)$ from (24) and the upper bound on $\bm{\widetilde{x}}_{i}$ by the truncation in (85). Therefore, for all $i\in[n]$, we have that
${\left\lVert\bm{\widetilde{x}}_{i}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{\widetilde{x}}_{i};1]\rangle\right\rVert}_{\psi_{2}}\overset{}{\leq}C{\left\lVert{\bm{x}}_{i}\right\rVert}_{\psi_{2}}\left|\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{\widetilde{x}}_{i};1]\rangle\right|\leq C\varsigma\sqrt{\log(end)},$ (91)
where the last inequality follows from (D.2). Therefore, we have by Hoeffding’s inequality and the union bound over $d$ entries that
$\mathbb{P}\left(\left|{\left\lVert\mathcal{M}_{2}\right\rVert}_{\infty}\right|\geq\epsilon\right)\leq 2d\exp\left\{\frac{-n\epsilon^{2}}{2C\varsigma^{2}\log(nd)}\right\}.$ (92)
The next statement follows by the union bound where we inflate the probability of error by $d$. Therefore, with $\epsilon=5\varsigma\log(nd)/\sqrt{n}$, we have that
$\mathbb{P}\left({\left\lVert\mathcal{M}_{2}\right\rVert}_{\infty}\geq C\varsigma\frac{\log(nd)}{\sqrt{n}}\right)\leq n^{-12}.$ (93)
For the last term, we consider the event $\bigcap\limits_{i=1}^{n}\mathcal{F}_{i}$. Under this event it holds with probability at least $1-n^{-11}$ that
${\left\lVert\frac{1}{n}\sum\limits_{i=1}^{n}z_{i}\bm{x}_{i}\right\rVert}_{\infty}\leq C\sigma_{z}\sqrt{\log(n)}{\left\lVert\frac{1}{n}\sum\limits_{i=1}^{n}\bm{x}_{i}\right\rVert}_{\infty}.$ (94)
The term on the right-hand side can be bounded by Hoeffding’s inequality and the union bound over $d$ entries as
$\mathbb{P}\left({\left\lVert\frac{1}{n}\sum\limits_{i=1}^{n}[\bm{x}_{i}]_{l}\right\rVert}_{\infty}\geq\epsilon\right)\leq 2d\exp\left\{\frac{-n\epsilon^{2}}{C}\right\},$
for some index $l\in[d]$. Finally, setting $\epsilon=(12\log(nd)/n)^{1/2}$, we get
$\mathbb{P}\left({\left\lVert\frac{1}{n}\sum\limits_{i=1}^{n}\bm{x}_{i}\right\rVert}_{\infty}\geq C\sqrt{\frac{12\log(nd)}{n}}\right)\leq 2n^{-12}.$ (95)
Combining (94) and (95), we get
$\mathbb{P}\left({\left\lVert\frac{1}{n}\sum\limits_{i=1}^{n}z_{i}\bm{x}_{i}\right\rVert}_{\infty}\geq C\sigma_{z}\frac{\log(nd)}{\sqrt{n}}\right)\leq 2n^{-11}.$ (96)
Finally, we combine (89),(93), and (96) to get
${\left\lVert\bm{m}_{1}-\bm{\widehat{m}}_{1}\right\rVert}_{\infty}\leq{\left\lVert\mathcal{M}_{1}\right\rVert}_{\infty}+{\left\lVert\mathcal{M}_{2}\right\rVert}_{\infty}+{\left\lVert\frac{1}{n}\sum\limits_{i=1}^{n}z_{i}\bm{x}_{i}\right\rVert}_{\infty}\leq C(\varsigma+\sigma_{z})\frac{\log(nd)}{\sqrt{n}},$ (97)
with probability at least $1-n^{-11}$. We proceed in a similar fashion and decompose the difference of the second moments as
$\widehat{\bm{M}}_{2}-\bm{M}_{2}$ $=\frac{1}{n}\sum_{i=1}^{n}\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d})-\mathbb{E}\left[\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d})\right]+\frac{1}{n}\sum_{i=1}^{n}z_{i}(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d})$
$=\underbrace{\frac{1}{n}\sum_{i=1}^{n}\left(\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}})-\mathbb{E}\left[\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}})\right]\right)}_{\mathcal{Q}_{1}}$
$\quad-\underbrace{\bm{I}_{d}\cdot\frac{1}{n}\sum_{i=1}^{n}\left(\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\widetilde{\bm{x}}_{i};1]\rangle-\mathbb{E}\left[\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\widetilde{\bm{x}}_{i};1]\rangle\right]\right)}_{\mathcal{Q}_{2}}+\frac{1}{n}\sum_{i=1}^{n}z_{i}(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d})$ (98)
To proceed, we define a generalization of sub-exponential and sub-Gaussian random variables as sub-Weibull random variables [Bakhshizadeh, 2023].
Definition D.1.
A random variable X is sub-Weibull of order $\alpha>0$, denoted as SW($\alpha$), if its sub-Weibull norm satisfies
${\left\lVert X\right\rVert}_{\psi_{\alpha^{-1}}}=\inf\left\{t:\exp\left(\left|\frac{X}{t}\right|^{\alpha^{-1}}\right)\leq 2\right\}<\infty.$ (99)
Note that sub-Gaussian random variables are SW($1/2$), and sub-exponential random variables are SW($1$). Next, we provide a useful lemma on algebra of sub-Weibull norms.
Lemma D.2.
Let X be SW($\alpha$) and Y be SW($\theta$) then we have that
${\left\lVert XY\right\rVert}_{\psi_{(\alpha+\theta)^{-1}}}\leq{\left\lVert X\right\rVert}_{\psi_{\alpha^{-1}}}\cdot{\left\lVert Y\right\rVert}_{\psi_{\theta^{-1}}}.$ (100)
Proof.
By the sub-Weibull assumptions we have that
$\mathbb{E}\left[\exp\left(\frac{X^{\alpha^{-1}}}{{\left\lVert X\right\rVert}_{\psi_{\alpha^{-}1}}}\right)\right]\leq 2,\quad\mathbb{E}\left[\exp\left(\frac{Y^{\theta^{-1}}}{{\left\lVert Y\right\rVert}_{\psi_{\theta^{-}1}}}\right)\right]\leq 2.$ (101)
Now we recall Young’s inequality as
$ab\leq\frac{a^{p}}{p}+\frac{b^{q}}{q},\quad\forall a,b\in\mathbb{R}^{+},\quad\forall p,q>1,\quad p^{-1}+q^{-1}=1.$ (102)
We observe that one valid pair for Young’s inequality is $(p,q)=\left(\frac{\alpha+\theta}{\alpha},\frac{\alpha+\theta}{\theta}\right)$. Therefore, we have that
$\mathbb{E}\left[\exp\left\{\frac{(XY)^{(\alpha+\theta)^{-1}}}{{\left\lVert X\right\rVert}_{\psi_{\alpha^{-}1}}{\left\lVert Y\right\rVert}_{\psi_{\theta^{-}1}}}\right\}\right]$
$\quad\leq\mathbb{E}\left[\exp\left\{\frac{\alpha}{\alpha+\theta}\cdot\frac{X^{\alpha^{-1}}}{{\left\lVert X\right\rVert}_{\psi_{\alpha^{-}1}}{\left\lVert Y\right\rVert}_{\psi_{\theta^{-}1}}}\right\}\cdot\exp\left\{\frac{\theta}{\alpha+\theta}\cdot\frac{Y^{\theta^{-1}}}{{\left\lVert X\right\rVert}_{\psi_{\alpha^{-}1}}{\left\lVert Y\right\rVert}_{\psi_{\theta^{-}1}}}\right\}\right]$
$\quad\leq\mathbb{E}\left[\frac{\alpha}{\alpha+\theta}\cdot\exp\left\{\frac{X^{\alpha^{-1}}}{{\left\lVert X\right\rVert}_{\psi_{\alpha^{-}1}}{\left\lVert Y\right\rVert}_{\psi_{\theta^{-}1}}}\right\}+\frac{\theta}{\alpha+\theta}\cdot\exp\left\{\frac{Y^{\theta^{-1}}}{{\left\lVert X\right\rVert}_{\psi_{\alpha^{-}1}}{\left\lVert Y\right\rVert}_{\psi_{\theta^{-}1}}}\right\}\right]$
$\quad\leq\mathbb{E}\left[\frac{\alpha}{\alpha+\theta}\cdot\exp\left\{\frac{X^{\alpha^{-1}}}{{\left\lVert X\right\rVert}_{\psi_{\alpha^{-}1}}\left({\left\lVert Y\right\rVert}_{\psi_{\theta^{-}1}}\vee 1\right)}\right\}+\frac{\theta}{\alpha+\theta}\cdot\exp\left\{\frac{Y^{\theta^{-1}}}{\left({\left\lVert X\right\rVert}_{\psi_{\alpha^{-}1}}\vee 1\right){\left\lVert Y\right\rVert}_{\psi_{\theta^{-}1}}}\right\}\right]$
$\quad\leq\frac{2\alpha}{\alpha+\theta}+\frac{2\theta}{\alpha+\theta}=2,$ (103)
where the first two inequalities follow from Young’s inequality which concludes the proof. ∎
We next present a very useful corollary for the case of product of sub-exponential and sub-Gaussian random variables.
Corollary D.3.
Let $\{X_{i}\}_{i=1}^{n_{x}}$ be sub-Gaussian random variables and $\{Y_{i}\}_{i=1}^{n_{y}}$ be sub-exponential random variables, then we have
${\left\lVert\prod\limits_{i=1}^{n_{x}}X_{i}\prod\limits_{j=1}^{n_{y}}Y_{j}\right\rVert}_{\psi_{({n_{x}}/{2}+n_{y})^{-1}}}\leq\prod\limits_{i=1}^{n_{x}}{\left\lVert X\right\rVert}_{\psi_{2}}\prod\limits_{j=1}^{n_{y}}{\left\lVert Y\right\rVert}_{\psi_{1}}$ (104)
Next, we show that each entry of $\mathcal{Q}_{1}$ is SW($3/2$). Let $l,m\in[d]$, then for all $i\in[n]$ we have
${\left\lVert\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle([\bm{x}_{i}]_{l}[\bm{x}_{i}]_{m})\right\rVert}_{\psi_{2/3}}$ $\leq{\left\lVert\max_{j\in[k]}\langle\bm{\theta}^{\star}_{j},[\bm{x}_{i};1]\rangle\right\rVert}_{\psi_{2}}{\left\lVert[\bm{x}_{i}]_{l}\right\rVert}_{\psi_{2}}{\left\lVert[\bm{x}_{i}]_{m}\right\rVert}_{\psi_{2}}$
$\leq C\sqrt{\log(d)}\max_{j\in[k]}\left({\left\lVert\bm{a}_{j}^{\star}\right\rVert}_{2}+|b_{j}^{\star}|\right){\left\lVert\bm{x}_{i}\right\rVert}^{2}_{\psi_{2}}\leq C\varsigma\sqrt{\log(d)},$ (105)
where the first inequality follows by Corollary D.3. We can now invoke the generalized version of Hoefdding’s and Bernstein’s inequality for sub-Weibull random variables. Using [Zhang and Wei, 2022, Proposition 3] with the union bound over $d^{2}$ entries, we have that
$\mathbb{P}\left({\left\lVert\mathcal{Q}_{1}\right\rVert}_{\infty}\geq C\varsigma\sqrt{\log(d)}\left(\sqrt{\frac{\epsilon}{n}}\vee\sqrt[3/2]{\frac{\epsilon}{n}}\right)\right)\leq d^{2}e^{-\epsilon},$ (106)
since the entries of $\mathcal{Q}_{1}$ are SW($3/2$). Choosing $\epsilon=12\log(nd)$ we get
$\mathbb{P}\left({\left\lVert\mathcal{Q}_{1}\right\rVert}_{\infty}\geq C\varsigma\sqrt{\log(d)}\left(\sqrt{\frac{\log(nd)}{n}}\vee\sqrt[3/2]{\frac{\log(nd)}{n}}\right)\right)\leq n^{-12}.$ (107)
Recalling (D.2), we have that the diagonal elements of $\mathcal{Q}_{2}$ are sub-Gaussian. Therefore, using Hoeffding’s inequality and the union bound of the $d$ diagonal entries we have
$\mathbb{P}\left({\left\lVert\mathcal{Q}_{2}\right\rVert}_{\infty}\geq\epsilon\right)\leq 2d\exp\left\{\frac{-n\epsilon^{2}}{2C\varsigma^{2}\log(nd)}\right\}.$ (108)
Choosing $\epsilon=5\varsigma\log(nd)/\sqrt{n}$ we get
$\mathbb{P}\left({\left\lVert\mathcal{Q}_{2}\right\rVert}_{\infty}\geq C\varsigma\frac{\log(nd)}{\sqrt{n}}\right)\leq n^{-12}.$ (109)
Next, for some $l,m\in[d]$, let the Kronecker delta $\delta_{lm}=1$ only when $l=m$ and 00 otherwise. Then we have for all $i\in[n]$ that
${\left\lVert z_{i}\left([\bm{x}_{i}]_{l}[\bm{x}_{i}]_{m}-\delta_{lm}\right)\right\rVert}_{\psi_{2/3}}\leq{\left\lVert z_{i}\right\rVert}_{\psi_{2}}{\left\lVert\bm{x}_{i}\right\rVert}^{2}_{\psi_{2}}\leq C\sigma_{z}.$ (110)
where the first inequality follows by Corollary D.3. Therefore, using [Zhang and Wei, 2022, Proposition 3] with the union bound over $d^{2}$ entries, we have that
$\mathbb{P}\left({\left\lVert\frac{1}{n}\sum_{i=1}^{n}z_{i}(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d})\right\rVert}_{\infty}\geq C\sigma_{z}\left(\sqrt{\frac{\log(nd)}{n}}\vee\sqrt[3/2]{\frac{\log(nd)}{n}}\right)\right)\leq n^{-12},$ (111)
Finally, we combine (107), (109), and (111) to get the second assertion in the lemma which concludes the proof.
D.3 Proof of Lemma B.3
We first derive a lower bound on
$\mathbb{P}\left(\bm{x}\in\mathcal{C}_{j}\cap\mathcal{C}_{j}^{\star}\right)$ $=\mathbb{P}\left(\bm{x}\in\mathcal{C}_{j}|\bm{x}\in\mathcal{C}_{j}^{\star}\right)\cdot\mathbb{P}\left(\bm{x}\in\mathcal{C}_{j}^{\star}\right)$
$=\left(1-\mathbb{P}\left(\bm{x}\not\in\mathcal{C}_{j}|\bm{x}\in\mathcal{C}_{j}^{\star}\right)\right)\cdot\mathbb{P}\left(\bm{x}\in\mathcal{C}_{j}^{\star}\right).$ (112)
Then, by the construction of $\{\mathcal{C}^{\star}_{j}\}_{j=1}^{k}$ in (12), we have
$\mathbb{P}\left(\bm{x}\not\in\mathcal{C}_{j}|\bm{x}\in\mathcal{C}_{j}^{\star}\right)=\frac{\mathbb{P}(\bm{x}\not\in\mathcal{C}_{j},\bm{x}\in\mathcal{C}_{j}^{\star})}{\mathbb{P}(\bm{x}\in\mathcal{C}_{j}^{\star})}$
[…]
$\leq\frac{1}{\pi_{j}^{\star}}\sum_{j^{\prime}\neq j}\mathbb{P}\left(\langle[\bm{x};1],\bm{v}_{j,j^{\prime}}^{\star}\rangle^{2}\leq\langle[\bm{x};1],\bm{v}_{j,j^{\prime}}-\bm{v}_{j,j^{\prime}}^{\star}\rangle^{2}\right),$
where the second inequality holds since $\bm{v}_{j,j^{\prime}}=\bm{\theta}_{j}-\bm{\theta}_{j^{\prime}}$ and $\bm{v}_{j,j^{\prime}}^{\star}=\bm{\theta}_{j}^{\star}-\bm{\theta}_{j^{\prime}}^{\star}$, and the last inequality follows from the fact that $ab\leq 0$ implies $|b|\leq|a-b|$ for $a,b\in\mathbb{R}$. Recall that $\bm{\theta}\in\mathcal{N}(\bm{\theta}^{\star},\Delta\rho)$ implies $\|\bm{v}_{j,j^{\prime}}-\bm{v}_{j,j^{\prime}}^{\star}\|_{2}\leq 2\rho\|(\bm{v}_{j,j^{\prime}}^{\star})_{1:d}\|_{2}$ due to [Kim and Lee, 2024, Lemma 7.4]. Furthermore, one can choose the absolute constant $R>0$ in (16) sufficiently small (but independent of $k$ and $d$) so that $2\rho\leq 0.1$. Then it follows that
$\mathbb{P}(\bm{x}\not\in\mathcal{C}_{j}\,|\,\bm{x}\in\mathcal{C}_{j}^{\star})$
$\overset{\mathrm{(i)}}{\leq}C\frac{k}{\pi_{j}^{\star}}\left(\frac{\|\bm{v}_{j,j^{\prime}}-\bm{v}_{j,j^{\prime}}^{\star}\|_{2}^{2}}{\|(\bm{v}_{j,j^{\prime}}^{\star})_{1:d}\|_{2}^{2}}\log\left(\frac{2\|(\bm{v}_{j,j^{\prime}}^{\star})_{1:d}\|_{2}}{\|\bm{v}_{j,j^{\prime}}-\bm{v}_{j,j^{\prime}}^{\star}\|_{2}}\right)\right)^{\zeta}$
$\overset{\mathrm{(ii)}}{\leq}C\frac{k}{\pi_{j}^{\star}}\left((2\rho)^{2}\log\left(\frac{1}{\rho}\right)\right)^{\zeta}$
$\overset{\mathrm{(iii)}}{\leq}C\frac{k}{\pi_{j}^{\star}}{\left(\frac{R^{2}\pi_{\min}^{2\zeta^{-1}(1+\zeta^{-1})}}{k^{2\zeta^{-1}}}\right)^{\zeta}}\overset{\mathrm{(iv)}}{\leq}C\frac{R^{2\zeta}\pi_{\min}^{1+2\zeta^{-1}}}{k}\overset{\mathrm{(v)}}{\leq}C\frac{R^{2\zeta}}{k^{2(1+\zeta^{-1})}}\triangleq\varrho,$ (113)
where (i) follows from [Kim and Lee, 2024, Lemma 7.5]; (ii) holds since $a\log^{1/2}(2/a)$ is monotone increasing for $a\in(0,1]$; (iii) follows from the fact that $a\leq\frac{b}{2}\log^{-1/2}(1/b)$ implies $a\log^{1/2}(2/a)\leq b$ for $b\in(0,0.1]$; (iv) holds since $\pi^{\star}_{j}\geq\pi_{\mathrm{min}}$ for all $j\in[k]$ by the defintion of $\pi_{\mathrm{min}}$ in (14); and (v) holds since $\pi_{\mathrm{min}}\leq\frac{1}{k}$. Once again $R>0$ can be made sufficiently small so that the right-hand side of (113) defined as $\varrho$ is arbitrarily small. Then plugging in this upper bound by (113) into (112) yields
$\pi_{j}\geq\mathbb{P}(\bm{x}\in\mathcal{C}_{j}\cap\mathcal{C}_{j}^{\star})\geq(1-\varrho)\cdot\pi_{j}^{\star}.$ (114)
Similarly, and by symmetry, we can write
$\mathbb{P}(\bm{x}\not\in\mathcal{C}_{j}\,|\,\bm{x}\in\mathcal{C}_{j}^{\star})$ $\leq\frac{R^{2\zeta}\pi_{\min}^{2(1+\zeta^{-1})}}{\pi_{j}k}\leq\frac{\pi_{j}^{\star}}{\pi_{j}}\cdot\frac{R^{2\zeta}\pi_{\min}^{1+2\zeta^{-1}}}{k}\leq\frac{\pi_{j}^{\star}}{\pi_{j}}\cdot\frac{R^{2\zeta}}{k^{2(1+\zeta^{-1})}}\leq\frac{\varrho}{1-\varrho},$ (115)
where the last inequality follows from the definition of $\varrho$ in (113) and the bound derived in (114). Finally, we can write the bound
$\pi_{j}^{\star}\geq\mathbb{P}(\bm{x}\in\mathcal{C}_{j}\cap\mathcal{C}_{j}^{\star})\geq\left(\frac{1-2\varrho}{1-\varrho}\right)\pi_{j}.$ (116)
Combining (114) and (116) provides the assertion in (75) thus concluding the proof.
D.4 Proof of Lemma B.4
Let $\mathcal{D}$ be a collection of subsets in $\mathbb{R}^{d}$. We denote the set of vectors whose entries are the indicator functions of $\mathcal{D}$ evaluated at samples $\{\bm{x}_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$ by
$\mathcal{H}(\mathcal{C},\{\bm{x}_{i}\}_{i=1}^{n})\triangleq\left\{\left(\mathbb{1}_{\{\bm{x}_{1}\in C\}},\ldots,\mathbb{1}_{\{\bm{x}_{n}\in C\}}\right):C\in\mathcal{D}\right\}.$ (117)
The Sauer-Shelah lemma (e.g. [Mohri et al., 2018, Section 3]) implies
$\sup\limits_{\bm{x}_{1},\ldots,\bm{x}_{n}}\left|\mathcal{H}(\mathcal{D},\{\bm{x}_{i}\}_{i=1}^{n})\right|\leq\left(\frac{en}{\operatorname{VC-dim}(\mathcal{D})}\right)^{\operatorname{VC-dim}(\mathcal{D})},$ (118)
where $\operatorname{VC-dim}(\mathcal{D})$ denotes the Vapnik-Chervonenkis dimension of $\mathcal{D}$. Recall that each elements of $\mathcal{P}_{k,d}(\varphi)$ is given as the intersection of $k$ jointly $s$-sparse halfspaces. Since the VC-dim of a single halfspace in $\mathbb{R}^{s}$ is $s+1$ [Csikos et al., 2018, Theorem A], we have
$\sup_{\bm{x}_{1},\dots,\bm{x}_{n}}\left|\mathcal{H}(\mathcal{P}_{k,d}(\varphi),\{\bm{x}_{i}\}_{i=1}^{n})\right|\leq\left(\frac{en}{s+1}\right)^{k(s+1)},\quad\forall\varphi\in\mathcal{Z}_{s}.$ (119)
Furthemore, since $\mathcal{P}_{k,d,s}=\underset{\varphi\in\mathcal{Z}_{s}}{\bigcup}\mathcal{P}_{k,d}(\varphi)$, it follows that
$\mathcal{H}(\mathcal{P}_{k,d,s},\{\bm{x}_{i}\}_{i=1}^{n})=\underset{\varphi\in\mathcal{Z}_{s}}{\bigcup}\mathcal{H}(\mathcal{P}_{k,d}(\varphi),\{\bm{x}_{i}\}_{i=1}^{n}).$ (120)
Therefore, the cardinality of $\mathcal{H}(\mathcal{P}_{k,d,s},\{\bm{x}_{i}\}_{i=1}^{n})$ satisfies
$\sup_{\bm{x}_{1},\dots,\bm{x}_{n}}\left|\mathcal{H}(\mathcal{P}_{k,d,s},\{\bm{x}_{i}\}_{i=1}^{n})\right|\leq\left(\frac{en}{s+1}\right)^{k(s+1)}\left(\frac{ed}{s}\right)^{s}.$ (121)
Then, to upper bound the term in (76), we use
$\sup\limits_{\begin{subarray}{c}C\in\mathcal{P}_{k,d,s}\\ \varphi\in\mathcal{Z}_{s}\end{subarray}}\left\|\bm{\widetilde{\Pi}}_{\varphi}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in C\}}\left(\bm{\xi}_{i}\bm{\xi}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\varphi}\right\|$
$\quad\leq\sup_{\bm{x}_{1}^{\prime},\ldots,\bm{x}_{n}^{\prime}}~{}\sup_{\{\omega_{i}\}_{i=1}^{n}\in\mathcal{H}(\mathcal{P}_{k,d,s},\{\bm{x}_{i}^{\prime}\}_{i=1}^{n})}\underbrace{\sup_{\varphi\in\mathcal{Z}_{s}}\left\|\bm{\widetilde{\Pi}}_{\varphi}\left(\frac{1}{n}\sum_{i=1}^{n}\omega_{i}\left(\bm{\xi}_{i}\bm{\xi}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\varphi}\right\|}_{f(\omega_{1},\dots,\omega_{n})}.$ (122)
Let $\{\omega_{i}\}_{i=1}^{n}\in\{0,1\}^{n}$ be arbitrarily fixed and $\alpha\triangleq\frac{1}{n}\sum_{i=1}^{n}\omega_{i}$. Then, via Lemma B.1, we obtain a tail bound on the sparsity-restricted spectral norm in the right-hand side of (D.4) given by
$\mathbb{P}\left(f(\omega_{1},\dots,\omega_{n})>\epsilon\right)\leq{\delta},$ (123)
if $n\geq C(\eta\vee 1)^{4}\epsilon^{-2}\alpha\left[s\log\left(\frac{d}{s}\right)+\log\left(\frac{1}{\delta}\right)\right]$. Trivially we have that $\alpha\leq 1$. Then, using the union bound, and inflating the probability of error $\delta$ by the worst-case cardinality in (121), we obtain that
$\mathbb{P}\left(\sup\limits_{\begin{subarray}{c}\bm{x}_{1}^{\prime},\ldots,\bm{x}_{n}^{\prime}\\ \{\omega_{i}\}_{i=1}^{n}\in\mathcal{H}(\mathcal{P}_{k,d,s},\{\bm{x}_{i}^{\prime}\}_{i=1}^{n})\end{subarray}}f(\omega_{1},\dots,\omega_{n})>\epsilon\right)\leq{\delta},$ (124)
if $n\geq C(\eta\vee 1)^{4}\epsilon^{-2}\left[sk\log\left(\frac{n}{s}\right)+s\log\left(\frac{d}{s}\right)+\log\left(\frac{1}{\delta}\right)\right]$ which concludes the proof.
D.5 Proof of Lemma A.1
This lemma is defined instate of the assumptions in 2.3. The proof of this lemma will simply involve invoking several auxiliary lemmas from Section B which hold with high probability. We will now prove that each of the statements in Lemma A.1 holds with probability at least $51-\delta/$ 5. Statement (47) follows from Lemma B.4 since (17) implies (77) with $\epsilon=k^{-3/2}\pi_{\mathrm{min}}^{2(1+\zeta^{-1})}$, and (48) holds from Lemma B.2 since (17) implies (74) with $\epsilon=k^{-3/2}\pi_{\mathrm{min}}^{2(1+\zeta^{-1})}$. Also, (49) follows from Lemma B.5 since $\bm{\theta}^{t}\in\mathcal{N}(\bm{\theta}^{\star},\Delta\rho)$ and (17) implies (80). Statement (50) follows directly from Lemma B.6. Finally, (51) holds directly from Lemma B.1 since (17) implies (71) with $\epsilon=k^{-2}\pi_{\mathrm{min}}^{2(1+\zeta^{-1})}$.